### v3

# Eka-Eval : A Comprehensive Framework for Indian Large Language Models

Abstract
The rapid advancement of Large Language Models (LLMs) has intensified the need for evaluation frameworks that address the requirements of linguistically diverse regions, such as India, and go beyond English-centric benchmarks. We introduce Eka-Eval, a unified evaluation framework that integrates over 35+ benchmarks (including 10 Indic benchmarks) across nine major evaluation categories. The Framework provides broader coverage than existing Indian language evaluation tools, offering 11 core capabilities through a modular architecture, seamless integration with Hugging Face and proprietary models, and plug-and-play usability. As the first end-to-end suite for scalable, multilingual LLM benchmarking, the framework combines extensive benchmarks, modular workflows, and dedicated support for low-resource Indian languages to enable inclusive assessment of LLM capabilities across diverse domains. We conducted extensive comparisons against five existing baselines, demonstrating that Eka-Eval achieves the highest participant ratings in four out of five categories.
[…]
Eka-Eval : A Comprehensive Framework for Indian Large Language Models
[…]
1 Introduction
Large Language Models (LLMs) have rapidly transformed natural language processing (NLP), enabling impressive generalization across diverse tasks including instruction following, reasoning, summarization, translation, and tool use. With the advent of general-purpose foundation models such as GPT-4 (Achiam et al., 2023), Claude (Anthropic, 2024), Gemini (Anil et al., 2023) and Llama-3 (Touvron et al., 2023), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems. Evaluation plays a critical role not only in measuring progress but in identifying capabilities, exposing limitations, and informing deployment strategies.
Figure 1: Architecture of Eka-Eval: A modular framework combining model inference, benchmark registry, evaluation engine, and result processing with support for caching and distributed inference.
In response, several evaluation frameworks have emerged, including HELM (Liang et al., 2022), EleutherAI Harness (Gao et al., 2021), lm-eval-harness (Gao et al., 2021), and OpenCompass (OpenCompass Contributors, 2023). However, existing evaluation frameworks are primarily centered on English or other high-resource languages and offer limited support for low-resource or multilingual settings (Watts et al., 2024). This lack of inclusivity significantly limits the effectiveness of these tools in linguistically diverse regions like India, which has 22 constitutionally recognized languages encompassing over a billion native speakers.
While notable benchmarks like IndicGLUE (Kakwani et al., 2020), IndicGenBench (Singh et al., 2024), and MILU (Verma et al., 2024) have addressed some gaps in Indic-language evaluation, they remain siloed efforts and lack integration into unified evaluation workflows — requiring users to perform manual configuration to use them in real-world LLM assessment pipelines.
Furthermore, popular frameworks such as FreeEval (Yu et al., 2024) and lm-eval-harness (Gao et al., 2021) require extensive configuration and engineering expertise, limiting their adoption by developers and researchers operating in low-resource environments. These challenges create a need for an actively maintained, community driven, multilingual, and task-diverse evaluation suite.
Framework Custom Datasets Custom Models Custom Prompting Long Context Tool Use Distributed Inference Visual Analysis Multilingual Indic Production Optimization Interactive CLI Quantization
lm-eval-harness
OpenCompass
HELM
OpenAI Evals
DeepEval
FreeEval
indic-eval
Eka-Eval
Table 1: Comparison of LLM Evaluation Frameworks across eleven key capabilities, showing the comprehensive coverage of Eka-Eval.
To address these limitations, we introduce Eka-Eval, a unified, extensible, and ready-to-use evaluation framework for LLMs, integrating over 35 benchmarks spanning both global (English) and Indic-language tasks. The overall architecture, presented in Figure 1, is designed to be modular, compatible with Hugging Face and proprietary models, it offers plug-and-play usability with minimal configuration overhead. To contextualize Eka-Eval’s capabilities, Table 1 compares leading LLM evaluation frameworks across eleven key capabilities, demonstrating its more comprehensive feature coverage relative to existing frameworks.
Eka-Eval covers nine major evaluation categories: (i) Code Generation and programming; (ii) Mathematics and logical reasoning; (iii) Reading comprehension; (iv) Commonsense reasoning; (v) World knowledge; (vi) Long-context understanding; (vii) General reasoning and knowledge; (viii) Tool use and API reasoning; and (viiii) Indic-specific NLP benchmarks including ARC-C (Clark et al., 2018), BoolQ (Clark et al., 2019), GSM8K (Cobbe et al., 2021), Math (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020a), SQuAD (Rajpurkar et al., 2018), APIBench (Patil et al., 2024), IndicGenBench (Singh et al., 2024).
Our key contributions are:
  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 35 benchmarks across nine major evaluation categories, providing comprehensive assessment of both global and Indic-language LLM capabilities.
  • We provide comprehensive support for most Indian languages and conduct thorough human evaluation across five existing frameworks on four benchmark tasks, demonstrating Eka-Eval’s effectiveness and reliability.
2 Related Work
The evaluation of LLMs has evolved from task-specific benchmarks to comprehensive, modular frameworks that support diverse capabilities and deployment settings. This section categorizes prior work into three major areas: general-purpose evaluation frameworks, specialized capability benchmarks, and multilingual/Indic evaluations. We conclude by highlighting how Eka-Eval integrates and advances these directions.
General-Purpose LLM Evaluation Frameworks: Early benchmark suites such as GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) laid the groundwork for multi-task evaluation, but were primarily English-centric and biased toward high-resource languages. Frameworks like HELM  (Liang et al., 2022) expanded evaluation to multidimensional axes—accuracy, calibration, robustness, fairness, and efficiency—applied uniformly across 30+ models. Similarly, BIG-Bench (Srivastava et al., 2022) curated over 200 tasks via crowdsourced contributions, highlighting model generality but with limited support for multilingualism or real-world deployment constraints. In parallel, tooling infrastructure also emerged. lm-eval-harness (Gao et al., 2021) provided reproducible few-shot evaluation pipelines, while OpenAI’s Evals (OpenAI Contributors, 2023) offered YAML-based structured evaluation for internal alignment research.
Recent frameworks emphasize modularity, extensibility, and accessibility. For instance, OpenCompass OpenCompass Contributors (2023) integrates distributed inference and real-time leaderboards, though it lacks Indic language support. FreeEval Yu et al. (2024) introduces meta-evaluation and contamination detection, but lacks robust multilingual and Indic language support. DeepEval DeepEval Contributors (2024) supports long-context and tool-use tasks but lacks flexible multilingual and low-resource customization, and its default metrics are often too rigid.
Specialized Capability Benchmarks: To evaluate emerging LLM capabilities, task-specific frameworks and datasets have been proposed. Tool-use and agentic behavior are assessed in ToolBench (Qin et al., 2023) and API-Bank (Li et al., 2023) respectively, which evaluate model interactions with real-world APIs.
Long-context reasoning is explored in InfiniteBench (Zhang et al., 2024) and RULER (Hsieh et al., 2024), which evaluate memory and coherence over 100K+ token sequences. Mathematical reasoning and code generation are evaluated using datasets like GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), respectively. While powerful, these benchmarks have limited scope and lack integration into unified evaluation pipelines.
Multilingual and Indic Language Evaluation: Multilingual evaluation has gained prominence through benchmark suites such as XTREME Hu et al. (2020), xP3 Muennighoff et al. (2022), and MEGA Ahuja et al. (2023), covering 40–70 languages. However, these benchmarks often rely on automatic translation, which may not reflect natural language usage or cultural nuances. FLORES Goyal et al. (2022) improved translation evaluation with high-quality parallel corpora for low-resource languages.
For Indian languages, IndicGLUE (Kakwani et al., 2020) and IndicXTREME (Kakwani et al., 2020) pioneered NLU benchmarking across major Indic languages. IndicGenBench (Singh et al., 2024) broadened the scope to generation tasks over 29 Indic languages, but remains a dataset collection rather than a complete evaluation framework, lacking system-level extensibility or plug-and-play usability. Frameworks such as indic-eval (IndicEval Contributors, 2024) wrap existing tools, like LightEval (LightEval, 2024), to support select Indic tasks. However, they offer limited extensibility and lack support for long-context tasks, tool-use evaluation, and custom prompting—features increasingly essential in real-world use cases (Chen et al., 2023).
Eka-Eval is a unified, extensible evaluation framework that integrates the breadth of multilingual and capability-centric benchmarks within a single, production-ready system. It supports a wide spectrum of evaluation settings, including both global and Indic benchmarks spanning 29+ Indian languages, designed for practical deployment and large-scale experimentation.
3 Capabilities of LLM Evaluation Frameworks
We identify eleven foundational capabilities that are critical for the design and continuous advancement of modern LLM evaluation frameworks. Several of these capabilities are described in earlier evaluation frameworks such as lm-eval-harness (Gao et al., 2021), OpenCompass (OpenCompass Contributors, 2023), FreeEval (Yu et al., 2024), and DeepEval (DeepEval Contributors, 2024). These eleven capabilities are:
(1) Custom Datasets:Support for loading and evaluating user-defined datasets beyond standard benchmark corpora.
(2) Custom Models: Compatibility with a wide range of models, including local checkpoints and API-hosted endpoints.
(3) Custom Prompting: Flexible, template-based prompting paradigms encompassing zero-shot, few-shot, and chain-of-thought configurations.
(4) Long Context: Ability to process and evaluate tasks involving extended input contexts (e.g., exceeding 4,000 tokens).
(5) Tool Use: Evaluating LLMs exhibiting agent-like behavior, including tool use, external API calls, and autonomous multi-step decision-making.
(6) Distributed Inference: Support for parallelized evaluation across multiple processes or compute nodes.
(7) Visual Analysis: Generation of interpretable visualizations, including bar charts, radar plots, and heatmaps, to facilitate comparative analysis.
(8) Multilingual (Indic): Native support for evaluation on benchmarks in Indic languages.
(9) Production Optimization: Implementation of runtime optimizations such as batching, caching, and prompt reuse to enhance evaluation efficiency.
(10) Interactive CLI: Availability of a command-line interface for interactive configuration of datasets, models, prompting strategies, and visualizations.
(11) Quantization: Compatibility with quantized model formats (e.g., 8-bit, 4-bit weights) to minimize memory requirements.
To contextualize the contributions of Eka-Eval, we benchmark it against established alternatives. Table 1 presents a comparative analysis of Eka-Eval and seven widely used LLM evaluation frameworks across the eleven capabilities outlined in this work. Eka-Eval demonstrates robust support across all key capabilities.
4 Eka-Eval
4.1 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances comprehensive benchmark coverage with practical usability. The system is designed around three core principles: modularity for easy extension and customization; accessibility for low-barrier adoption across diverse research environments; and comprehensiveness supports a wide range of capabilities, including underserved areas such as long-context reasoning and tool use.
4.2 System Architecture
The framework follows a layered architecture consisting of four primary components:
4.2.1 Evaluation Engine
This component orchestrates all evaluation workflows:
Task Scheduler: Manages task scheduling, prompt formatting, and result aggregation across distributed inference setups. The scheduler implements intelligent work distribution as demonstrated in main_orchestrator() - dynamically assigning evaluation tasks to available workers based on resource constraints and model requirements.
Batch Optimizer: Implements intelligent batching strategies and supports various quantization schemes to optimize memory usage and inference speed. As seen in the PIQA evaluation implementation, the optimizer automatically adjusts generation_batch_size parameters to maximize throughput while preventing out-of-memory errors.
Distributed Coordinator: Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library. The coordinator launches multiple worker_process instances to handle independent evaluation tasks, enabling parallel execution across different benchmarks and model configurations.
4.2.2 Benchmark Registry
Provides a unified interface for managing datasets:
Dataset Manager: The BenchmarkRegistry class handles diverse dataset formats and sources, abstracting the complexities of different evaluation protocols. The manager supports datasets from HuggingFace Hub, local files, and custom data formats through standardized interfaces.
4.2.3 Model Interface Layer
Abstracts access to different local and API-based model backends:
Local Model Loader: Initializes transformer-based checkpoints with automatic device allocation and quantization.
API Client Manager: Manages proprietary endpoints through dedicated clients (OpenAIClient, GeminiClient, ClaudeClient) that extend BaseAPIClient, providing unified request handling with rate limiting and authentication.
Interactive Selection Interface: Implements get_model_selection_interface() for dynamic model discovery and selection, supporting local model paths and API configurations.
Resource Manager: Ensures efficient memory management through explicit cleanup functions, preventing resource leaks during repeated evaluation runs.
Framework Setup & Config Navigation Result Export Indic Support Extensibility
[…]
4.2.4 Results Processing System
Handles comprehensive output management through three secondary components:
Metrics Calculator: Computes evaluation metrics using HuggingFace’s evaluate library (e.g., accuracy, BLEU, F1-score, exact match, Pass@1). It also implements robust error handling for edge cases and missing data; for example, when a model returns “The answer is probably B”, regex-based extraction retrieves the label; if that fails, a default score is assigned. Code completions are sandboxed with timeout control to ensure safe execution and error logging.
Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps and radar plots (including support for cross-model performance comparisons).
Export Manager: Handles result export in multiple formats including JSON and CSV. The manager maintains evaluation metadata including model parameters, benchmark versions, execution timestamps, and system configurations.
4.3 Comprehensive Benchmark Coverage
Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 35+ benchmarks (See Appendix A). These categories include:
  1. 1.
    Code Generation and Programming: Programming abilities are assessed using HumanEval (Chen et al., 2021), MBPP  (Austin et al., 2021), HumanEval+ (Liu et al., 2023), EvalPlus (Liu et al., 2023), with Pass@1 accuracy metrics.
  2. 2.
    Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K (Cobbe et al., 2021) for grade school math, MATH (Hendrycks et al., 2021) for competition-level problems, and ARC-C (Clark et al., 2018) for scientific reasoning.
  3. 3.
    Reading Comprehension: Text understanding capabilities are evaluated using SQuAD Rajpurkar et al. (2018), QuAC Choi et al. (2018), and BoolQ Clark et al. (2019) with F1 scores and exact match metrics.
  4. 4.
    Commonsense Reasoning: We incorporate PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), ARC-C (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), CommonSenseQA (Talmor et al., 2018), and OpenBookQA (Mihaylov et al., 2018) for comprehensive commonsense evaluation.
  5. 5.
    World Knowledge: Factual knowledge is tested through TriviaQA (5-shot) (Joshi et al., 2017) and NaturalQuestions (5-shot) (Kwiatkowski et al., 2019) with accuracy metrics.
  6. 6.
    Long-Context Understanding: For extended context reasoning, we include ZeroSCROLLS (Shaham et al., 2023) with ROUGE (Lin, 2004) and F1 scores, Needle-in-a-Haystack (Wang et al., 2024a) for retrieval accuracy, and InfiniteBench (Zhang et al., 2024) for task-specific long-context evaluation.
  7. 7.
    General Reasoning and Knowledge: For foundational capabilities, we include MMLU (Hendrycks et al., 2020a) and MMLU-Pro (Wang et al., 2024b) for multitask language understanding, IFEval (Zhou et al., 2023) for instruction following, BBH (3-shot) (Suzgun et al., 2022) for challenging reasoning tasks, and AGI-Eval (3-5 shot) (Zhong et al., 2023) for general intelligence assessment.
  8. 8.
    Tool Use and API Reasoning: Practical capabilities are assessed through API-Bank Li et al. (2023) for API call accuracy and ROUGE-L and API-Bench Patil et al. (2024) for API recommendation accuracy.
  9. 9.
    Multilingual and Indic Language Support: A distinguishing feature of Eka-Eval is its dedicated support for Indian languages, addressing a critical gap in existing evaluation frameworks. The system includes benchmarks covering the majority of Indian languages. For multilingual evaluation, it supports popular Indic language benchmarks including MMLU-IN (Hendrycks et al., 2020b), TriviaQA-IN (Joshi et al., 2017), MILU (Verma et al., 2024), GSM8K-IN (Cobbe et al., 2021), BoolQ-IN (Clark et al., 2019), ARC-C-IN (Clark et al., 2018), Flores-IN (Goyal et al., 2022) with BLEU (Papineni et al., 2002) and ChrF metrics (Popović, 2015), XQuAD-IN (Artetxe et al., 2019) and XorQA-IN (Asai et al., 2021) with F1 and exact match scores.
    The multilingual architecture includes language-specific prompts, culturally appropriate protocols, and specialized tokenization for Indic scripts, ensuring fair evaluation across languages while staying consistent with global benchmarks.
4.4 Extensibility and Customization
The framework is designed with extensibility as a first-class concern. New benchmarks can be integrated through a simple plugin architecture that requires minimal boilerplate code. The system supports custom evaluation metrics, prompt templates, and post-processing pipelines, enabling users to adapt the framework to their specific requirements.
Configuration is managed through a hierarchical JSON-based system that enables users to define evaluation suites ranging from quick smoke tests to comprehensive benchmark runs. It also supports parameter sweeps, facilitating systematic exploration of prompt variations, few-shot examples, and model hyperparameters.
5 Experiments
To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval.
5.1 Evaluation Procedure
Three human participants independently evaluated each framework by running the Sarvam-1B Sarvam (2024) model on four diverse benchmarks: WinoGrande, PIQA, ARC-C and FLORES. Participants installed each framework in a standardized environment, integrated the target model, and ran the recommended benchmark workflows. They then rated the framework on five criteria using Likert scale ranging from 1 (Very Poor) to 5 (Excellent):
EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks. Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility. Participants were encouraged to consult official documentation but not to seek external technical support beyond publicly available resources.
Participants assessed each framework across five evaluation criteria:
  1. 1.
    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and execute an initial benchmark run.
  2. 2.
    Ease of Navigation: Intuitiveness of navigation, benchmark selection, and configuration, including CLI clarity, documentation quality, and ease of discovering options.
  3. 3.
    Result Reporting and Export: Clarity and accessibility of evaluation outputs, with options to export results (e.g., JSON, CSV) and create visualizations such as bar charts or heatmaps.
  4. 4.
    Indic Language Support: Support for Indic and multilingual datasets such as ARC-IN and FLORES. Evaluators assessed the availability of prebuilt resources.
  5. 5.
    Extensibility: Ease of customizing the framework to add new prompt templates, models, datasets, or evaluation metrics.
[…]
6 Conclusion and Future Work
In this work, we introduced Eka-Eval, a unified and extensible framework designed to streamline the evaluation of LLMs across 35+ diverse benchmarks and 11 key capabilities, spanning nine major evaluation categories-including support for Indic languages. Through a combination of system design, implementation, and user-centered evaluation, Eka-Eval demonstrated high usability and practicality, achieving the highest participant ratings in four out of five evaluation criteria. These results highlight its effectiveness in enabling reproducible, scalable, and inclusive evaluation workflows.
We plan to expand the framework to over 100 tasks, with a focus on underrepresented Indic languages. Future work will categorize benchmarks by language diversity, task difficulty, and domains such as law, healthcare, and governance. We plan to add dynamic task calibration for context length, ambiguity, reasoning complexity, and India-specific knowledge tasks. Priorities include benchmarks for bias detection, hallucination analysis, privacy risks, domain-specific assessments, and adversarial human review. Finally, we will support phase-wise evaluation across pretraining, fine-tuning, and deployment to provide a comprehensive view of model behavior. These enhancements will be included in Eka-Eval v2.0, strengthening its role as the evaluation standard for Indic and other LLMs.
7 Ethics Statement
This research uses publicly available data without personally identifiable information. All datasets and models comply with their terms of use. The work is intended for academic research. Potential misuse or unintended amplification of biases should be carefully considered before deployment.
Appendix A Appendix
A.1 Indic benchmark configuration
As demonstrated in Figure 2, a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Indic languages. It illustrates how task parameters, templates, and dataset references are modularly specified in Eka-Eval.
"ARC-Challenge-Indic": {
  "description": "Zero-shot evaluation
    across 11 Indic languages",
  "evaluation_function":
    "indic.arc_c_in.evaluate_arc_c_in",
  "task_args": {
    "dataset_name":
      "sarvamai/arc-challenge-indic",
    "target_languages": ["bn", "en",
      "gu", "hi", "kn", "ml", "mr",
      "or", "pa", "ta", "te"],
    "dataset_split": "validation",
    "num_few_shot": 0,
    "max_new_tokens": 10,
    "generation_batch_size": 8,
    "prompt_template_name_zeroshot":
      "arc_c_in_0shot",
    "prompt_template_name_fewshot":
      "arc_c_in_5shot",
    "prompt_file_benchmark_key":
      "arc_c_in",
    "prompt_file_category": "indic"
Figure 2: ARC-Challenge-Indic benchmark configuration example
A.2 CLI Demonstration
The interactive CLI of the Eka-Eval framework is shown below, which guides users through model selection and evaluation setup. Simplifying benchmarking workflows, it is accessible to both researchers and developers.
Figure 3: Available benchmarks groups of Eka-Eval framework.
As per Figure 3, users are prompted to select high-level task groups (e.g., Reading Comprehension) during CLI setup. This enables fine-grained benchmarking organization and streamlined selection.
Figure 4: Model selection in the Eka-Eval framework.
Eka-Eval supports local HuggingFace models and API-based models like Sarvam, Gemma, OpenAI, Claude, and Gemini. Users interactively select model source and configuration through CLI. (see Figure 4)
Figure 5: Subtask selection within a task group.
After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain. (see Figure 5)
Figure 6: Consolidated evaluation results table.
The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV. (see Figure 6)
Figure 7: Interactive visualisation setup in Eka-Eval.
As per Figure 7) framework allows users to generate multiple types of visualizations: bar charts, heatmaps, radar plots that are based on completed evaluations.
Figure 8: Bar chart visualisation
Figure 8 shows performance breakdown across sub-tasks like BoolQ, SQuAD, and QuAC. It provides intuitive insight into strengths and weaknesses of the model.
A.2.1 Prompt Template System
A critical component of Eka-Eval is its sophisticated prompt management system, which handles diverse evaluation paradigms and languages. The framework implements a flexible template system demonstrated through PIQA benchmark prompt 9:
  "piqa_generation": {
    "template": "Choose the most appropriate
      solution (0 or 1) to achieve the goal:
[…]
    "description": "Generation-based PIQA prompt"
  "piqa_5shot_generation": {
    "template_prefix": "Choose the most
      appropriate solution (0 or 1)...",
    "few_shot_example_template":
[…]
    "template_suffix": "Question: {goal}
[…]
    "description": "Few-shot generation prompt"
  "default_few_shot_examples_piqa": [
      "goal": "To remove a stain from clothing",
      "sol1": "Apply cold water immediately...",
      "sol2": "Set the clothing on fire...",
      "answer_label": "0"
Figure 9: PIQA prompt templates supporting multiple evaluation paradigms
The Prompt template system as shown in Figure 9 supports zero-shot, few-shot, and chain-of-thought prompting strategies, ensuring consistency across evaluation modes and languages. Users can customize prompt strategies and easily configure them in the benchmark configuration file, as shown in Figure 2.
Appendix B Limitations
While Eka-Eval supports a wide range of benchmarks and model backends, it currently lacks a graphical user interface, relying instead on CLI-based workflows. The framework does not yet support vLLMs and has limited support for detailed error analysis or explainability. Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
Appendix C Acknowledgements
The authors express their gratitude to Ritika Iyer, Nikhil Mishra, and Rajat Kumar Thakur for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results. We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Shailesh Panda.

### v4

@@
-Eka-Eval : A Comprehensive Framework for Indian Large Language Models
+Eka-Eval [Uncaptioned image] : An Evaluation Framework for Low-Resource Multilingual Large Language Models
 Abstract
-The rapid advancement of Large Language Models (LLMs) has intensified the need for evaluation frameworks that address the requirements of linguistically diverse regions, such as India, and go beyond English-centric benchmarks.
-We introduce Eka-Eval, a unified evaluation framework that integrates over 35+ benchmarks (including 10 Indic benchmarks) across nine major evaluation categories.
-The Framework provides broader coverage than existing Indian language evaluation tools, offering 11 core capabilities through a modular architecture, seamless integration with Hugging Face and proprietary models, and plug-and-play usability.
-As the first end-to-end suite for scalable, multilingual LLM benchmarking, the framework combines extensive benchmarks, modular workflows, and dedicated support for low-resource Indian languages to enable inclusive assessment of LLM capabilities across diverse domains.
-We conducted extensive comparisons against five existing baselines, demonstrating that Eka-Eval achieves the highest participant ratings in four out of five categories.
-[…]
-[…]
-[…]
-Eka-Eval : A Comprehensive Framework for Indian Large Language Models
+The rapid evolution of Large Language Models’ has underscored the need for evaluation frameworks that are globally applicable, flexible, and modular, and that support a wide range of tasks, model types, and linguistic settings.
+We introduce Eka-Eval, a unified, end-to-end framework that combines a zero-code web interface and an interactive CLI to ensure broad accessibility.
+It integrates 50+ multilingual benchmarks across nine evaluation categories, supports local and proprietary models, and provides 11 core capabilities through a modular, plug-and-play architecture.
+Designed for scalable, multilingual evaluation with support for low-resource multilingual languages, Eka-Eval is, to the best of our knowledge, the first suite to offer comprehensive coverage in a single platform.
+Comparisons against five existing baselines indicate improvements of at least 2x better on key usability measures, with the highest user satisfaction, faster setup times, and consistent benchmark reproducibility.
+Website Eka-Eval
+Demo bit.ly/Eka-Eval
+Code github.com/lingo-iitgn/eka-eval
+Eka-Eval [Uncaptioned image] : An Evaluation Framework for Low-Resource Multilingual Large Language Models
 […]
 1 Introduction
-Large Language Models (LLMs) have rapidly transformed natural language processing (NLP), enabling impressive generalization across diverse tasks including instruction following, reasoning, summarization, translation, and tool use.
-With the advent of general-purpose foundation models such as GPT-4 (Achiam et al., 2023), Claude (Anthropic, 2024), Gemini (Anil et al., 2023) and Llama-3 (Touvron et al., 2023), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems.
+Large Language Models’ (LLMs) growing capabilities continue to reshape NLP, enabling impressive generalization across diverse tasks, including instruction following, reasoning, summarization, translation, and tool use.
+With the advent of general-purpose foundation models such as GPT-4 (achiam2023gpt), Claude (anthropic2024claude35sonnet), Gemini (team2023gemini), Deepseek (liu2024deepseek), and Llama-3 (touvron2023llama), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems.
 Evaluation plays a critical role not only in measuring progress but in identifying capabilities, exposing limitations, and informing deployment strategies.
 Figure 1: Architecture of Eka-Eval: A modular framework combining model inference, benchmark registry, evaluation engine, and result processing with support for caching and distributed inference.
-In response, several evaluation frameworks have emerged, including HELM (Liang et al., 2022), EleutherAI Harness (Gao et al., 2021), lm-eval-harness (Gao et al., 2021), and OpenCompass (OpenCompass Contributors, 2023).
-However, existing evaluation frameworks are primarily centered on English or other high-resource languages and offer limited support for low-resource or multilingual settings (Watts et al., 2024).
-This lack of inclusivity significantly limits the effectiveness of these tools in linguistically diverse regions like India, which has 22 constitutionally recognized languages encompassing over a billion native speakers.
-While notable benchmarks like IndicGLUE (Kakwani et al., 2020), IndicGenBench (Singh et al., 2024), and MILU (Verma et al., 2024) have addressed some gaps in Indic-language evaluation, they remain siloed efforts and lack integration into unified evaluation workflows — requiring users to perform manual configuration to use them in real-world LLM assessment pipelines.
-Furthermore, popular frameworks such as FreeEval (Yu et al., 2024) and lm-eval-harness (Gao et al., 2021) require extensive configuration and engineering expertise, limiting their adoption by developers and researchers operating in low-resource environments.
-These challenges create a need for an actively maintained, community driven, multilingual, and task-diverse evaluation suite.
-Framework Custom Datasets Custom Models Custom Prompting Long Context Tool Use Distributed Inference Visual Analysis Multilingual Indic Production Optimization Interactive CLI Quantization
+Core Flexibility Advanced Capabilities Usability & Specialization
+Framework
+Custom
+Data
+Custom
+Models
+Custom
+Prompt
+Quant-
+ization
+Long
+Context
+Tool
+Use
+Dist.
+Infer.
+Visual
+Analysis
+Zero-Code
+UI
+Interact.
+CLI
+Low-Resource
+Multilingual
 lm-eval-harness
@@
 Eka-Eval
-Table 1: Comparison of LLM Evaluation Frameworks across eleven key capabilities, showing the comprehensive coverage of Eka-Eval.
-To address these limitations, we introduce Eka-Eval, a unified, extensible, and ready-to-use evaluation framework for LLMs, integrating over 35 benchmarks spanning both global (English) and Indic-language tasks.
-The overall architecture, presented in Figure 1, is designed to be modular, compatible with Hugging Face and proprietary models, it offers plug-and-play usability with minimal configuration overhead.
-To contextualize Eka-Eval’s capabilities, Table 1 compares leading LLM evaluation frameworks across eleven key capabilities, demonstrating its more comprehensive feature coverage relative to existing frameworks.
-Eka-Eval covers nine major evaluation categories: (i) Code Generation and programming; (ii) Mathematics and logical reasoning; (iii) Reading comprehension; (iv) Commonsense reasoning; (v) World knowledge; (vi) Long-context understanding; (vii) General reasoning and knowledge; (viii) Tool use and API reasoning; and (viiii) Indic-specific NLP benchmarks including ARC-C (Clark et al., 2018), BoolQ (Clark et al., 2019), GSM8K (Cobbe et al., 2021), Math (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020a), SQuAD (Rajpurkar et al., 2018), APIBench (Patil et al., 2024), IndicGenBench (Singh et al., 2024).
-Our key contributions are:
-  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 35 benchmarks across nine major evaluation categories, providing comprehensive assessment of both global and Indic-language LLM capabilities.
-  • We provide comprehensive support for most Indian languages and conduct thorough human evaluation across five existing frameworks on four benchmark tasks, demonstrating Eka-Eval’s effectiveness and reliability.
-2 Related Work
-The evaluation of LLMs has evolved from task-specific benchmarks to comprehensive, modular frameworks that support diverse capabilities and deployment settings.
-This section categorizes prior work into three major areas: general-purpose evaluation frameworks, specialized capability benchmarks, and multilingual/Indic evaluations.
-We conclude by highlighting how Eka-Eval integrates and advances these directions.
-General-Purpose LLM Evaluation Frameworks: Early benchmark suites such as GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) laid the groundwork for multi-task evaluation, but were primarily English-centric and biased toward high-resource languages.
-Frameworks like HELM  (Liang et al., 2022) expanded evaluation to multidimensional axes—accuracy, calibration, robustness, fairness, and efficiency—applied uniformly across 30+ models.
-Similarly, BIG-Bench (Srivastava et al., 2022) curated over 200 tasks via crowdsourced contributions, highlighting model generality but with limited support for multilingualism or real-world deployment constraints.
-In parallel, tooling infrastructure also emerged.
-lm-eval-harness (Gao et al., 2021) provided reproducible few-shot evaluation pipelines, while OpenAI’s Evals (OpenAI Contributors, 2023) offered YAML-based structured evaluation for internal alignment research.
-Recent frameworks emphasize modularity, extensibility, and accessibility.
-For instance, OpenCompass OpenCompass Contributors (2023) integrates distributed inference and real-time leaderboards, though it lacks Indic language support.
-FreeEval Yu et al.
-(2024) introduces meta-evaluation and contamination detection, but lacks robust multilingual and Indic language support.
-DeepEval DeepEval Contributors (2024) supports long-context and tool-use tasks but lacks flexible multilingual and low-resource customization, and its default metrics are often too rigid.
-Specialized Capability Benchmarks: To evaluate emerging LLM capabilities, task-specific frameworks and datasets have been proposed.
-Tool-use and agentic behavior are assessed in ToolBench (Qin et al., 2023) and API-Bank (Li et al., 2023) respectively, which evaluate model interactions with real-world APIs.
-Long-context reasoning is explored in InfiniteBench (Zhang et al., 2024) and RULER (Hsieh et al., 2024), which evaluate memory and coherence over 100K+ token sequences.
-Mathematical reasoning and code generation are evaluated using datasets like GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), respectively.
-While powerful, these benchmarks have limited scope and lack integration into unified evaluation pipelines.
-Multilingual and Indic Language Evaluation: Multilingual evaluation has gained prominence through benchmark suites such as XTREME Hu et al.
-(2020), xP3 Muennighoff et al.
-(2022), and MEGA Ahuja et al.
-(2023), covering 40–70 languages.
-However, these benchmarks often rely on automatic translation, which may not reflect natural language usage or cultural nuances.
-FLORES Goyal et al.
-(2022) improved translation evaluation with high-quality parallel corpora for low-resource languages.
-For Indian languages, IndicGLUE (Kakwani et al., 2020) and IndicXTREME (Kakwani et al., 2020) pioneered NLU benchmarking across major Indic languages.
-IndicGenBench (Singh et al., 2024) broadened the scope to generation tasks over 29 Indic languages, but remains a dataset collection rather than a complete evaluation framework, lacking system-level extensibility or plug-and-play usability.
-Frameworks such as indic-eval (IndicEval Contributors, 2024) wrap existing tools, like LightEval (LightEval, 2024), to support select Indic tasks.
-However, they offer limited extensibility and lack support for long-context tasks, tool-use evaluation, and custom prompting—features increasingly essential in real-world use cases (Chen et al., 2023).
-Eka-Eval is a unified, extensible evaluation framework that integrates the breadth of multilingual and capability-centric benchmarks within a single, production-ready system.
-It supports a wide spectrum of evaluation settings, including both global and Indic benchmarks spanning 29+ Indian languages, designed for practical deployment and large-scale experimentation.
+Table 1: Feature comparison with state-of-the-art frameworks.
+Takeaway: Eka-Eval seamlessly integrates an accessible dual interface (Zero-Code UI + Interactive CLI) and cutting-edge features (Long Context, Tool Use) with Low-Resource Multilingual benchmarks.
+An Alarming Need: Existing evaluation frameworks for LMs, such as HELM (liang2022helm), EleutherAI Harness (gao2021framework), lm-eval-harness (gao2021framework), and OpenCompass (wu2023opencompass), predominantly target high-resource languages and rely on command-line interfaces with limited benchmark coverage (watts2024pariksha).
+These frameworks create significant barriers for non-technical users who require graphical interfaces and researchers working with low-resource and multilingual settings.
+The challenge is especially pronounced in linguistically diverse regions globally, where numerous languages serve large populations; yet, existing evaluation frameworks lack comprehensive multilingual support and culturally grounded benchmarks, limiting rigorous assessment of language model capabilities across diverse linguistic settings (pomerenke2025ai; xuan-etal-2025-mmlu; vayani2025all).
+Motivation: Existing evaluation frameworks suffer from several limitations such as (1) absence of user-friendly interfaces, requiring extensive command-line expertise, broken-pipelines and manual configuration (gao2021framework; liang2022holistic), (2) restricted flexibility in benchmark selection and task customization (wang2018glue; srivastava2022beyond), (3) insufficient task diversity, with fragmented support for advanced capabilities like long-context reasoning and tool use (qin2023toolllm; zhang2024bench), and (4) limited multilingual coverage, particularly for low-resource Indian languages (hu2020xtreme; ahuja2023mega).
+To address this, we propose Eka-Eval, a uniquely scalable and extensible evaluation framework that unifies multilingual benchmarks under flexible task configurations.
+It enables seamless model integration and multi-platform accessibility, establishing a new benchmark for low-resource multilingual LLM evaluation frameworks.
+Contributions: The key contributions are:
+  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 50 multilingual benchmarks across nine major evaluation categories.
+  • We introduce a first-ever Zero-Code UI for benchmark evaluation in diverse and low-resource multilingual settings, along with visualisations, leaderboard, and an interactive CLI, making large-scale evaluation accessible to both non-technical and technical users.
+  • Through a comprehensive human evaluation involving eleven participants comparing five existing frameworks, showing that Eka-Eval achieves the highest participant ratings across all six features, performing at least 2× better than existing frameworks in key usability and effectiveness.
+2 Relevant Work
+General-Purpose Frameworks: Early multi-task suites such as GLUE (wang2018glue) and SuperGLUE (wang2019superglue) established foundational standards for evaluation.
+More comprehensive frameworks, including HELM (liang2022helm), BIG-Bench (srivastava2022beyond), lm-eval-harness (gao2021framework), and OpenAI Evals (openai2023evals) broaden task coverage and introduce extensible evaluation pipelines.
+yet typically require substantial effort and provide limited multilingual capabilities.
+Modular ecosystems such as OpenCompass (wu2023opencompass) and FreeEval (yu2024freeeval) offer distributed inference and advanced features, but still lack unified evaluation and zero-code accessibility.
+Specialized Capability Benchmarks: Several domain-specific benchmarks target advanced LLM abilities, such as tool use in ToolBench (qin2023toolbench), agentic behavior in API-Bank (li2023api) and InfiniteBench (zhang2024bench), long-context reasoning in RULER (hsieh2024ruler), and mathematical and coding competencies in GSM8K (cobbe2021training) and HumanEval (chen2021evaluating).
+While powerful, these benchmarks are fragmented across independent repositories, requiring users to mainitain multiple codebases to conduct a thorough evaluation.
+Multilingual Evaluation: Multilingual suites such as XTREME (hu2020xtreme) and MEGA (ahuja2023mega) cover 30+ languages but often rely on machine-translated content that lacks cultural grounding.
+Recent efforts of MCEval (Huang2025MCEvalAD) (13 cultures, 39K dynamic instances), GlotEval (Luo2025GlotEvalAT) (language-agnostic, culturally adaptive prompts), BLEnD (myung2024blend) (16 cultures, 52K native QA pairs), CulturalBench (chiu2024culturalbench) (45 global regions), and ALM-Bench (vayani2025all) (100 languages with cultural images) address this gap.
+However, these benchmarks still face insufficient data for low-resource languages, cultural biases favor dominant dialects, and evaluation metrics struggle to capture cultural appropriateness.
+Eka-Eval mitigates these gaps by unifying 50+ diverse benchmarks within a single modular workflow.
+It offers dual accessibility (zero-code UI + interactive CLI), advanced capabilities (long-context, tool-use, distributed inference), and multilingual language support for comprehensive evaluation.
 3 Capabilities of LLM Evaluation Frameworks
-We identify eleven foundational capabilities that are critical for the design and continuous advancement of modern LLM evaluation frameworks.
-Several of these capabilities are described in earlier evaluation frameworks such as lm-eval-harness (Gao et al., 2021), OpenCompass (OpenCompass Contributors, 2023), FreeEval (Yu et al., 2024), and DeepEval (DeepEval Contributors, 2024).
-These eleven capabilities are:
-(1) Custom Datasets:Support for loading and evaluating user-defined datasets beyond standard benchmark corpora.
-(2) Custom Models: Compatibility with a wide range of models, including local checkpoints and API-hosted endpoints.
-(3) Custom Prompting: Flexible, template-based prompting paradigms encompassing zero-shot, few-shot, and chain-of-thought configurations.
-(4) Long Context: Ability to process and evaluate tasks involving extended input contexts (e.g., exceeding 4,000 tokens).
-(5) Tool Use: Evaluating LLMs exhibiting agent-like behavior, including tool use, external API calls, and autonomous multi-step decision-making.
-(6) Distributed Inference: Support for parallelized evaluation across multiple processes or compute nodes.
-(7) Visual Analysis: Generation of interpretable visualizations, including bar charts, radar plots, and heatmaps, to facilitate comparative analysis.
-(8) Multilingual (Indic): Native support for evaluation on benchmarks in Indic languages.
-(9) Production Optimization: Implementation of runtime optimizations such as batching, caching, and prompt reuse to enhance evaluation efficiency.
-(10) Interactive CLI: Availability of a command-line interface for interactive configuration of datasets, models, prompting strategies, and visualizations.
-(11) Quantization: Compatibility with quantized model formats (e.g., 8-bit, 4-bit weights) to minimize memory requirements.
-To contextualize the contributions of Eka-Eval, we benchmark it against established alternatives.
-Table 1 presents a comparative analysis of Eka-Eval and seven widely used LLM evaluation frameworks across the eleven capabilities outlined in this work.
-Eka-Eval demonstrates robust support across all key capabilities.
+We identify eleven foundational capabilities critical for modern LLM evaluation frameworks, informed by established systems like lm-eval-harness (gao2021framework), OpenCompass (wu2023opencompass), FreeEval (yu2024freeeval), and DeepEval (deepeval).
+These capabilities are organized into three categories:
+A. Core Flexibility:
+(1) Custom Datasets: Support for user-defined evaluation data beyond standard benchmarks.
+(2) Custom Models: Compatibility with local checkpoints and API-hosted endpoints.
+(3) Custom Prompting: Flexible templates supporting zero-shot, few-shot, and chain-of-thought configurations.
+(4) Quantization: Support for memory-efficient quantized formats (8-bit, 4-bit weights).
+B. Advanced Capabilities:
+(5) Long Context: Processing extended inputs exceeding 4,000 tokens.
+(6) Tool Use: Evaluating agent-like behavior including API calls and multi-step reasoning.
+(7) Distributed Inference: Parallelized evaluation across multiple compute nodes.
+(8) Visual Analysis: Generation of interpretable visualizations (bar charts, radar plots, heatmaps, etc.).
+C. Usability & Specialization:
+(9) Zero-Code UI: Web-based interface enabling evaluation without programming.
+(10) Interactive CLI: Command-line interface for configuring datasets, models, and strategies.
+(11) Low-Resource Multilingual: Comprehensive multilingual benchmarks featuring language-specific tokenization.
+Table 1 presents a feature-level comparison demonstrating that Eka-Eval is the only framework offering complete coverage across technical, linguistic, and usability dimensions.
+While lm-eval-harness, OpenCompass, and HELM provide strong technical capabilities, they require substantial expertise and suffer from usability challenges.
+indic-eval, though tailored to Indic languages, supports only seven benchmarks and lacks broader functionality.
+Critically, existing frameworks lack zero-code web interfaces, restricting accessibility to users with programming expertise.
+Eka-Eval uniquely combines comprehensive evaluation capabilities with an intuitive zero-code UI, significantly lowering barriers to rigorous multilingual assessment.
 4 Eka-Eval
 4.1 Design and Implementation
-Eka-Eval is architected as a modular, extensible evaluation framework that balances comprehensive benchmark coverage with practical usability.
-The system is designed around three core principles: modularity for easy extension and customization; accessibility for low-barrier adoption across diverse research environments; and comprehensiveness supports a wide range of capabilities, including underserved areas such as long-context reasoning and tool use.
+Eka-Eval is architected as a modular, extensible evaluation framework that balances practical usability with comprehensive benchmark coverage.
+It provides a zero-code, web-based interface that allows users to run benchmarks, adjust parameters, and visualize results through interactive analytics and leaderboards, while a CLI enables detailed configuration and large-scale evaluations.
+The framework supports flexible model selection for both local and API-based models and is built around three core principles: modularity for seamless extension and customization; accessibility to support diverse environments, including low-resource language settings; and comprehensiveness to cover a wide range of capabilities and multilingual benchmarks, with particular attention to underserved areas such as long-context reasoning and tool use.
 4.2 System Architecture
-The framework follows a layered architecture consisting of four primary components:
+The framework has a layered architecture with four main components (illustrated in Figure 1):
 4.2.1 Evaluation Engine
-This component orchestrates all evaluation workflows:
+Manages all evaluation workflows:
 Task Scheduler: Manages task scheduling, prompt formatting, and result aggregation across distributed inference setups.
-The scheduler implements intelligent work distribution as demonstrated in main_orchestrator() - dynamically assigning evaluation tasks to available workers based on resource constraints and model requirements.
+The scheduler implements intelligent work distribution, as shown in main_orchestrator(), by dynamically assigning evaluation tasks to available workers based on resource constraints and model requirements.
 Batch Optimizer: Implements intelligent batching strategies and supports various quantization schemes to optimize memory usage and inference speed.
 As seen in the PIQA evaluation implementation, the optimizer automatically adjusts generation_batch_size parameters to maximize throughput while preventing out-of-memory errors.
-Distributed Coordinator: Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library.
-The coordinator launches multiple worker_process instances to handle independent evaluation tasks, enabling parallel execution across different benchmarks and model configurations.
+Distributed Coordinator: Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library, launching multiple worker_process instances to execute independent evaluation tasks in parallel across benchmarks and model configurations.
 4.2.2 Benchmark Registry
@@
 Dataset Manager: The BenchmarkRegistry class handles diverse dataset formats and sources, abstracting the complexities of different evaluation protocols.
-The manager supports datasets from HuggingFace Hub, local files, and custom data formats through standardized interfaces.
+The manager supports datasets from HuggingFace Hub, local files, and custom API-models through standardized interfaces.
 4.2.3 Model Interface Layer
-Abstracts access to different local and API-based model backends:
+Access to different local and API-based models:
 Local Model Loader: Initializes transformer-based checkpoints with automatic device allocation and quantization.
 API Client Manager: Manages proprietary endpoints through dedicated clients (OpenAIClient, GeminiClient, ClaudeClient) that extend BaseAPIClient, providing unified request handling with rate limiting and authentication.
-Interactive Selection Interface: Implements get_model_selection_interface() for dynamic model discovery and selection, supporting local model paths and API configurations.
-Resource Manager: Ensures efficient memory management through explicit cleanup functions, preventing resource leaks during repeated evaluation runs.
-Framework Setup & Config Navigation Result Export Indic Support Extensibility
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Interactive Selection Interface: Implements get_model_selection_interface() for dynamic model discovery and selection, supporting local model paths and API configurations(See Appendix C).
+Resource Manager: Ensures efficient memory management via cleanup functions, preventing resource leaks during repeated evaluation runs.
+Framework Setup & Config Navigation UI Result Export Extensibility Multilingual Support
+lm-eval-harness 3.73 $\pm$ 0.72 3.91 $\pm$ 0.66 1.00 $\pm$ 0.00 3.64 $\pm$ 0.70 4.18 $\pm$ 1.00 1.64 $\pm$ 0.50
+OpenCompass 2.18 $\pm$ 0.93 2.36 $\pm$ 0.67 3.09 $\pm$ 1.36 3.09 $\pm$ 0.93 2.80 $\pm$ 1.06 2.09 $\pm$ 0.71
+HELM 1.91 $\pm$ 0.71 2.64 $\pm$ 1.05 2.36 $\pm$ 1.22 2.55 $\pm$ 1.42 2.09 $\pm$ 0.93 1.36 $\pm$ 0.50
+indic-eval 2.55 $\pm$ 0.88 3.18 $\pm$ 0.67 1.00 $\pm$ 0.00 2.55 $\pm$ 1.00 2.55 $\pm$ 0.88 4.55 $\pm$ 0.53
+FreeEval 2.55 $\pm$ 1.24 2.45 $\pm$ 0.71 2.73 $\pm$ 1.01 2.64 $\pm$ 1.24 2.36 $\pm$ 0.87 1.45 $\pm$ 0.53
+Eka-Eval 4.27 $\pm$ 0.93 4.55 $\pm$ 0.73 4.64 $\pm$ 0.53 4.55 $\pm$ 0.50 4.64 $\pm$ 0.50 4.73 $\pm$ 0.50
+Table 2: Average participant ratings of the evaluation frameworks by eleven participants (mean $\pm$ standard deviation; Likert scale: 1–5, 1 = poor and 5 = excellent).
+Takeaway: Eka-Eval achieves the highest ratings across all functionalities, with participants consistently perceiving it as superior across the six assessed capabilities.
+   Framework    Time Taken
+   lm-eval-harness    $22\pm 7.58$
+   OpenCompass    $36\pm 9.83$
+   HELM    $58\pm 5.70$
+   IndicEval    $32\pm 4.14$
+   FreeEval    $35\pm 7.07$
+   Eka-Eval    $\mathbf{11\pm 3.18}$
+Table 3: Comparison of time taken for installation and configuration across six frameworks reported across five successful runs on GPU.
+Takeaway: Eka-Eval records the lowest setup time and exhibits consistently low variance.
 4.2.4 Results Processing System
 Handles comprehensive output management through three secondary components:
-Metrics Calculator: Computes evaluation metrics using HuggingFace’s evaluate library (e.g., accuracy, BLEU, F1-score, exact match, Pass@1).
-It also implements robust error handling for edge cases and missing data; for example, when a model returns “The answer is probably B”, regex-based extraction retrieves the label; if that fails, a default score is assigned.
-Code completions are sandboxed with timeout control to ensure safe execution and error logging.
-Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps and radar plots (including support for cross-model performance comparisons).
-Export Manager: Handles result export in multiple formats including JSON and CSV.
+Metrics Calculator: Employs HuggingFace’s industry-standard evaluate library to compute essential metrics such as accuracy, BLEU, F1-score, exact match, and Pass@1, ensuring consistency with widely adopted benchmarks and standard evaluation practices across diverse LLM tasks.
+Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps, and radar plots (including support for cross-model comparisons).
+Export Manager: Handles result export in multiple formats, including JSON and CSV.
 The manager maintains evaluation metadata including model parameters, benchmark versions, execution timestamps, and system configurations.
+4.2.5 Interface and Deployment Layer
+Handles complex evaluation logic with ease of use through a streamlined full-stack user interface (See Appendix §B and Figure 4).
+Full-Stack Architecture: A decoupled React frontend and FastAPI backend, served through Nginx, provide a stable and scalable deployment layer.
+Real-Time Telemetry: WebSocket-based streaming delivers live inference logs and GPU status directly to the UI, eliminating the need for CLI monitoring.
+LLM-Powered Diagnostics: An LLM (LLaMA-3.3 70B (grattafiori2024llama)) interprets evaluation outputs and generates summaries of model performance and failure patterns, revealing additional insights.
+Interactive Dashboard: A “Zero-Code” interface with visual comparisons and leaderboards enables users to run and analyze evaluations without technical overhead.
 4.3 Comprehensive Benchmark Coverage
-Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 35+ benchmarks (See Appendix A). These categories include:
+Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 50+ benchmarks (See Appendix A).
+These categories include:
   1. 1.
-    Code Generation and Programming: Programming abilities are assessed using HumanEval (Chen et al., 2021), MBPP  (Austin et al., 2021), HumanEval+ (Liu et al., 2023), EvalPlus (Liu et al., 2023), with Pass@1 accuracy metrics.
+    Code Generation and Programming: Programming abilities are assessed using HumanEval (chen2021evaluating), MBPP (austin2021program), HumanEval+ (liu2023your), and MBPP EvalPlus (liu2023your) with Pass@1 metrics.
   2. 2.
-    Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K (Cobbe et al., 2021) for grade school math, MATH (Hendrycks et al., 2021) for competition-level problems, and ARC-C (Clark et al., 2018) for scientific reasoning.
+    Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K (cobbe2021training), MATH (hendrycks2021measuring), and ARC-Challenge (clark2018think).
   3. 3.
-    Reading Comprehension: Text understanding capabilities are evaluated using SQuAD Rajpurkar et al.
-(2018), QuAC Choi et al.
-(2018), and BoolQ Clark et al.
-(2019) with F1 scores and exact match metrics.
+    Reading Comprehension: Text understanding is evaluated using SQuAD (rajpurkar2018know), QuAC (choi2018quac), and BoolQ (clark2019boolq) with F1 and exact match metrics.
   4. 4.
-    Commonsense Reasoning: We incorporate PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), ARC-C (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), CommonSenseQA (Talmor et al., 2018), and OpenBookQA (Mihaylov et al., 2018) for comprehensive commonsense evaluation.
+    Commonsense Reasoning: We incorporate PIQA (bisk2020piqa), SIQA (sap2019social), HellaSwag (zellers2019hellaswag), ARC-Easy/Challenge (clark2018think), WinoGrande (sakaguchi2021winogrande), CommonSenseQA (talmor2018commonsenseqa), and OpenBookQA (mihaylov2018can).
   5. 5.
-    World Knowledge: Factual knowledge is tested through TriviaQA (5-shot) (Joshi et al., 2017) and NaturalQuestions (5-shot) (Kwiatkowski et al., 2019) with accuracy metrics.
+    World Knowledge: Factual knowledge is tested through TriviaQA (joshi2017triviaqa) and NaturalQuestions (kwiatkowski2019natural).
   6. 6.
-    Long-Context Understanding: For extended context reasoning, we include ZeroSCROLLS (Shaham et al., 2023) with ROUGE (Lin, 2004) and F1 scores, Needle-in-a-Haystack (Wang et al., 2024a) for retrieval accuracy, and InfiniteBench (Zhang et al., 2024) for task-specific long-context evaluation.
+    Long-Context Understanding: For extended context, we include ZeroSCROLLS (shaham2023zeroscrolls), Needle-in-a-Haystack (wang2024multimodal), and InfiniteBench (zhang2024bench).
   7. 7.
-    General Reasoning and Knowledge: For foundational capabilities, we include MMLU (Hendrycks et al., 2020a) and MMLU-Pro (Wang et al., 2024b) for multitask language understanding, IFEval (Zhou et al., 2023) for instruction following, BBH (3-shot) (Suzgun et al., 2022) for challenging reasoning tasks, and AGI-Eval (3-5 shot) (Zhong et al., 2023) for general intelligence assessment.
+    General Reasoning: Foundational capabilities are tested via MMLU (hendrycks2021mmlu), MMLU-Pro (wang2024mmlu), IFEval (zhou2023instruction), BBH (suzgun2022challenging), and AGI-Eval (zhong2023agieval).
   8. 8.
-    Tool Use and API Reasoning: Practical capabilities are assessed through API-Bank Li et al.
-(2023) for API call accuracy and ROUGE-L and API-Bench Patil et al.
-(2024) for API recommendation accuracy.
+    Tool Use and API Reasoning: Practical capabilities are assessed through API-Bank (li2023api) and API-Bench (patil2023api).
   9. 9.
-    Multilingual and Indic Language Support: A distinguishing feature of Eka-Eval is its dedicated support for Indian languages, addressing a critical gap in existing evaluation frameworks.
-The system includes benchmarks covering the majority of Indian languages.
-For multilingual evaluation, it supports popular Indic language benchmarks including MMLU-IN (Hendrycks et al., 2020b), TriviaQA-IN (Joshi et al., 2017), MILU (Verma et al., 2024), GSM8K-IN (Cobbe et al., 2021), BoolQ-IN (Clark et al., 2019), ARC-C-IN (Clark et al., 2018), Flores-IN (Goyal et al., 2022) with BLEU (Papineni et al., 2002) and ChrF metrics (Popović, 2015), XQuAD-IN (Artetxe et al., 2019) and XorQA-IN (Asai et al., 2021) with F1 and exact match scores.
-    The multilingual architecture includes language-specific prompts, culturally appropriate protocols, and specialized tokenization for Indic scripts, ensuring fair evaluation across languages while staying consistent with global benchmarks.
+    Multilingual Language Support: Eka-Eval features the largest unified Multilingual suite (20 benchmarks), including: (i) Knowledge: IndicMMLU-Pro (sankalp2025indicmmlupro), MMLU-IN (hendrycks2020measuring), TriviaQA-IN (joshi2017triviaqa), MILU (verma2024milu).
+(ii) Reasoning: HellaSwag-IN (zellers2019hellaswag), ARC-C IN (clark2018think), IndicCOPA (kakwani2020indicnlpsuite), GSM8K-IN (cobbe2021training).
+(iii) Reading & QA: BoolQ-IN (clark2019boolq), XQuAD-IN (artetxe2019cross), XorQA-IN (asai2021xor), Indic-QA (singh2025indicqabenchmarkmultilingual).
+(iv) Generation (NLG): Flores-IN (goyal-etal-2022-flores), IndicParaphrase, IndicWikiBio, IndicQuestionGeneration, IndicSentenceSummarization, IndicHeadlineGeneration (singh2024indicgenbench).
+(v) NLU: IndicNER, IndicSentiment, IndicGLUE (kakwani2020indicnlpsuite).
+   Framework    Winogrande    PIQA    ARC-IN    MMLU-IN
+   lm-eval-harness    65.8 $\pm$ 6.0    74.8 $\pm$ 3.5    –    –
+   OpenCompass    61.1 $\pm$ 5.0    73.3 $\pm$ 4.3    –    –
+   indic-eval    65.8 $\pm$ 1.9    78.7 $\pm$ 1.4    28.1 $\pm$ 2.3    35.0 $\pm$ 2.1
+   FreeEval    60.0 $\pm$ 4.5    54.5 $\pm$ 3.2    –    –
+   Eka-Eval    66.5 $\pm$ 2.5    77.5 $\pm$ 1.5    31.5 $\pm$ 3.5    33.5 $\pm$ 0.5
+Table 4: Benchmark Score Parity.
+Comparison of benchmark scores (mean $\pm$ standard deviation).
+‘–’ represents those benchmarks are not supported for the frameworks.
+Takeaway: Eka-Eval reproduces verified consistent scores with established frameworks and closely aligns with their actual benchmark results.
 4.4 Extensibility and Customization
-The framework is designed with extensibility as a first-class concern.
-New benchmarks can be integrated through a simple plugin architecture that requires minimal boilerplate code.
-The system supports custom evaluation metrics, prompt templates, and post-processing pipelines, enabling users to adapt the framework to their specific requirements.
-Configuration is managed through a hierarchical JSON-based system that enables users to define evaluation suites ranging from quick smoke tests to comprehensive benchmark runs.
-It also supports parameter sweeps, facilitating systematic exploration of prompt variations, few-shot examples, and model hyperparameters.
+Eka-Eval emphasizes extensibility through a low-code plugin architecture that enables easy integration of new benchmarks and custom metrics with minimal modifications, managed via a hierarchical JSON configuration system that supports complex setups such as parameter sweeps and prompt variations.
+The framework provides a zero-code UI that enables real-time, interactive adjustment of inference parameters (e.g., temperature, batch size) and evaluation strategies, eliminating the need for users to manage configuration files directly.
 5 Experiments
 To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval.
-5.1 Evaluation Procedure
-Three human participants independently evaluated each framework by running the Sarvam-1B Sarvam (2024) model on four diverse benchmarks: WinoGrande, PIQA, ARC-C and FLORES.
-Participants installed each framework in a standardized environment, integrated the target model, and ran the recommended benchmark workflows.
-They then rated the framework on five criteria using Likert scale ranging from 1 (Very Poor) to 5 (Excellent):
+Evaluation Procedure
+Eleven graduate-level computer science researchers with demonstrated proficiency in Git/GitHub evaluated each framework in a single-blind setup.
+Participants followed a standardized evaluation protocol in a controlled environment.
+They installed each framework, integrated HuggingFace models - including gemma-2-2b (team2024gemma) and sarvam-1B sarvamai_sarvam1_2024, and executed the recommended workflows on four diverse benchmarks: WinoGrande, PIQA, ARC-C-IN and MMLU-IN.
+They then rated the framework on five criteria using a Likert scale ranging from 1 to 5:
 EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks.
-Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility.
-Participants were encouraged to consult official documentation but not to seek external technical support beyond publicly available resources.
-Participants assessed each framework across five evaluation criteria:
+Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility, with explicit dos and don’ts specified for all participants.
+They were permitted to consult official documentation but were prohibited from seeking external technical support beyond publicly available framework resources.
+Participants assessed each framework across six evaluation criteria:
   1. 1.
-    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and execute an initial benchmark run.
+    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and run an initial benchmark.
   2. 2.
@@
   3. 3.
+    Zero-Code UI Availability: Presence and usability of GUI-based tools for visualization, benchmark execution, and result analysis, eliminating the need for command-line interaction.
+  4. 4.
     Result Reporting and Export: Clarity and accessibility of evaluation outputs, with options to export results (e.g., JSON, CSV) and create visualizations such as bar charts or heatmaps.
-  4. 4.
-    Indic Language Support: Support for Indic and multilingual datasets such as ARC-IN and FLORES. Evaluators assessed the availability of prebuilt resources.
   5. 5.
     Extensibility: Ease of customizing the framework to add new prompt templates, models, datasets, or evaluation metrics.
-[…]
+  6. 6.
+    Multilingual Language Support: Support for multilingual benchmarks such as ARC-IN, FLORES, and similar. Evaluators assessed the availability of prebuilt resources.
+Table 2 presents average participant ratings across six evaluation criteria using a 1-5 Likert scale (1= poor or absence of functionality, 5 =excellent or flexible capability).
+Eka-Eval achieved the highest ratings across all criteria, establishing superior usability and capability.
+Moreover, Table 3 summarizes the average setup times across six frameworks, which varied widely due to dependency issues and unstable codebases in several baselines.
+Nine of the eleven participants reported that HELM suffered from poor documentation and a broken pipeline, while OpenCompass required manual code edits and indic-eval experienced installation difficulties.
+Eka-Eval achieved the shortest setup time.
+Finally, Table 4 presents reproduced benchmark results on gemma-2-2b, showing how they align with broader framework differences and highlighting Eka-Eval’s consistency across global-multilingual benchmarks.
+HELM was excluded from this comparison, as it neither supports these benchmarks nor produced functional results despite repeated attempts.
 6 Conclusion and Future Work
-In this work, we introduced Eka-Eval, a unified and extensible framework designed to streamline the evaluation of LLMs across 35+ diverse benchmarks and 11 key capabilities, spanning nine major evaluation categories-including support for Indic languages.
-Through a combination of system design, implementation, and user-centered evaluation, Eka-Eval demonstrated high usability and practicality, achieving the highest participant ratings in four out of five evaluation criteria.
-These results highlight its effectiveness in enabling reproducible, scalable, and inclusive evaluation workflows.
-We plan to expand the framework to over 100 tasks, with a focus on underrepresented Indic languages.
-Future work will categorize benchmarks by language diversity, task difficulty, and domains such as law, healthcare, and governance.
-We plan to add dynamic task calibration for context length, ambiguity, reasoning complexity, and India-specific knowledge tasks.
-Priorities include benchmarks for bias detection, hallucination analysis, privacy risks, domain-specific assessments, and adversarial human review.
-Finally, we will support phase-wise evaluation across pretraining, fine-tuning, and deployment to provide a comprehensive view of model behavior.
-These enhancements will be included in Eka-Eval v2.0, strengthening its role as the evaluation standard for Indic and other LLMs.
+We present Eka-Eval, an open-source unified evaluation framework designed to overcome persistent challenges in accessibility, reproducibility, and multilingual coverage for language model assessment.
+Our study with eleven participants demonstrates substantial usability improvements, achieving lowest setup time and the highest satisfaction ratings.
+Eka-Eval enables an interactive UI, rigorous, reproducible evaluation without manual integration overhead.
+The current release is an ongoing work, with 30 stars and 3 forks on GitHub.
+Future development will extend the framework to more than 100 tasks, specialized domains, and advanced diagnostic capabilities, establishing Eka-Eval 2.0 as a comprehensive and equitable standard for multilingual language model evaluation.
+7 Limitations
+While Eka-Eval supports a wide range of benchmarks and model backends, it does not yet support multi-modal LLMs.
+Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
 8 Ethics Statement
@@
 Potential misuse or unintended amplification of biases should be carefully considered before deployment.
+9 Acknowledgements
+The authors express their gratitude to Himanshu Beniwal for helping with writing the manuscript and to Birudugadda Srivibhav, Sailesh Panda, Gautham Bharati, Indrayudh Mandal, Krudant Randai, Sahil Gawande, Tirth Bhatt, Shruti Bhat, Mann agarwal, Himanshu Sharma and Khushbu Bijawat for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results.
+We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Mahavir Patil.
 Appendix A Appendix
-A.1 Indic benchmark configuration
-As demonstrated in Figure 2, a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Indic languages.
+A.1 Multilingual benchmark configuration
+As demonstrated in Figure 2, a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Multilingual languages.
 It illustrates how task parameters, templates, and dataset references are modularly specified in Eka-Eval.
@@
 Figure 2: ARC-Challenge-Indic benchmark configuration example
-A.2 CLI Demonstration
-The interactive CLI of the Eka-Eval framework is shown below, which guides users through model selection and evaluation setup.
-Simplifying benchmarking workflows, it is accessible to both researchers and developers.
-Figure 3: Available benchmarks groups of Eka-Eval framework.
-As per Figure 3, users are prompted to select high-level task groups (e.g., Reading Comprehension) during CLI setup. This enables fine-grained benchmarking organization and streamlined selection.
-Figure 4: Model selection in the Eka-Eval framework.
-Eka-Eval supports local HuggingFace models and API-based models like Sarvam, Gemma, OpenAI, Claude, and Gemini. Users interactively select model source and configuration through CLI. (see Figure 4)
-Figure 5: Subtask selection within a task group.
-After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain. (see Figure 5)
-Figure 6: Consolidated evaluation results table.
-The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV. (see Figure 6)
-Figure 7: Interactive visualisation setup in Eka-Eval.
-As per Figure 7) framework allows users to generate multiple types of visualizations: bar charts, heatmaps, radar plots that are based on completed evaluations.
-Figure 8: Bar chart visualisation
-Figure 8 shows performance breakdown across sub-tasks like BoolQ, SQuAD, and QuAC. It provides intuitive insight into strengths and weaknesses of the model.
-A.2.1 Prompt Template System
+A.2 Prompt Template System
 A critical component of Eka-Eval is its sophisticated prompt management system, which handles diverse evaluation paradigms and languages.
-The framework implements a flexible template system demonstrated through PIQA benchmark prompt 9:
+The framework implements a flexible template system demonstrated through PIQA benchmark prompt 3:
   "piqa_generation": {
@@
     "template_prefix": "Choose the most
-      appropriate solution (0 or 1)...",
     "few_shot_example_template":
@@
 Users can customize prompt strategies and easily configure them in the benchmark configuration file, as shown in Figure 2.
-Appendix B Limitations
-While Eka-Eval supports a wide range of benchmarks and model backends, it currently lacks a graphical user interface, relying instead on CLI-based workflows.
-The framework does not yet support vLLMs and has limited support for detailed error analysis or explainability.
-Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
-Appendix C Acknowledgements
-The authors express their gratitude to Ritika Iyer, Nikhil Mishra, and Rajat Kumar Thakur for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results.
-We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Shailesh Panda.
+Figure 4: High-level architectural overview of the Eka-Eval framework.
+Appendix B Zero-code Web Interface
+To democratize access for LLM evaluation, Eka-Eval includes a comprehensive "Zero-code" web interface.
+This GUI mirrors the capabilities of the CLI serving as a frontend layer as seen in (Figure 4) while adding advanced features for prompt configuration, resource management, and AI-driven analysis making it accessible for both non-technical and technical users.
+Figure 5: High-level benchmark selection dashboard categorized by domain (e.g., Math, Code Generation, Indic Benchmarks).
+The Benchmark Selection Dashboard
+(Figure 5) allows users to build their own evaluation suite by simply toggling and selecting the benchmarks available and also adding custom benchmarks.
+Figure 6: Advanced configuration panel for adjusting inference parameters (temperature, batch size) and managing GPU resources.
+For advanced control over the evaluation process, Configuration Panel (Figure 6) offers adjustable settings such as temperature, top-p and batch size.It also includes a built-in GPU resource manager, allowing users to allocate specific or multi CUDA for evaluation.
+Figure 7: The prompt customization interface, allowing users to modify templates and injection points for specific tasks.
+A key advantage of the web interface is the Prompt Customization Engine (Figure 7).
+It lets users visually edit system prompts, few-shot templates, and variable fields, removing the need to manually modify JSON configuration files.
+Figure 8: Post-evaluation summary showing overall progress, specific benchmark scores, and options to export results or view detailed analysis.
+The Evaluation Dashboard
+(Figure 8) shows real time progress during inference including a live console and, once finished, provides a summary of the model’s performance with options to download logs or explore detailed analysis.
+Figure 9: AI diagonosis dashboard providing automated textual insights into model strengths and weaknesses based on performance.
+AI Diagnosis Dashboard
+uses LLama 3.3-70b to analyze the evaluation logs, generating summaries of the model’s observed strengths (e.g., "Strong logical coherence") and weaknesses (e.g., "Hallucination in low-resource languages") as example shown in (Figure 9).
+Figure 10: AI diagnosis of the wrong result of benchmark selected
+Users can also inspect specific failure cases (Figure 10).
+Selecting a failed benchmark item shows the prompt, the model’s wrong answer, and an AI-generated explanation of why the error likely occurred, making it easy to spot and fix issues.
+Figure 11: Model Leaderboard featuring multi-dimensional visualizations (radar and bar charts) for comparative analysis.
+Finally, the Model Leaderboard (Figure 11) provides a central place to compare models.
+It gathers results from all local runs and visualizes performance differences with radar charts and bar graphs, helping users choose the best model.
+Appendix C Interactive CLI
+C.1 CLI Demonstration
+The interactive CLI of the Eka-Eval framework is shown below, which guides users through model selection and evaluation setup.
+Simplifying benchmarking workflows, it is accessible to both researchers and developers.
+Figure 12: Available benchmark groups of the Eka-Eval framework.
+As per Figure 12, users are prompted to select high-level task groups (e.g., Reading Comprehension) during CLI setup. This enables fine-grained benchmarking organization and streamlined selection.
+Figure 13: Model selection in the Eka-Eval framework.
+Figure 14: Interactive visualisation setup in Eka-Eval.
+Figure 15: Subtask selection within a task group.
+Figure 16: Bar chart visualisation
+After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain (see Figure 15).
+Figure 17: Consolidated evaluation results table.
+The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV (see Figure 17).
+Figure 16 shows performance breakdown across sub-tasks like BoolQ, SQuAD, and QuAC. It provides intuitive insight into strengths and weaknesses of the model.
+Appendix D Computation Requirement and Budget
+For computational infrastructure, experiments were carried out on four NVIDIA Tesla V100 32 GB GPUs, with an estimated cost of \$5,431.20 per month based on Google Cloud Platform (GCP)  [^1] https://cloud.google.com/products/calculator Calculator pricing.

### v5

@@
 Abstract
-The rapid evolution of Large Language Models’ has underscored the need for evaluation frameworks that are globally applicable, flexible, and modular, and that support a wide range of tasks, model types, and linguistic settings.
+The rapid evolution of Large Language Models has underscored the need for evaluation frameworks that are globally applicable, flexible, and modular, and that support a wide range of tasks, model types, and linguistic settings.
 We introduce Eka-Eval, a unified, end-to-end framework that combines a zero-code web interface and an interactive CLI to ensure broad accessibility.
-It integrates 50+ multilingual benchmarks across nine evaluation categories, supports local and proprietary models, and provides 11 core capabilities through a modular, plug-and-play architecture.
+It integrates 55+ diverse benchmarks across nine evaluation categories, supports local and proprietary models, and provides 11 core capabilities through a modular, plug-and-play architecture.
 Designed for scalable, multilingual evaluation with support for low-resource multilingual languages, Eka-Eval is, to the best of our knowledge, the first suite to offer comprehensive coverage in a single platform.
@@
 1 Introduction
-Large Language Models’ (LLMs) growing capabilities continue to reshape NLP, enabling impressive generalization across diverse tasks, including instruction following, reasoning, summarization, translation, and tool use.
+Large Language Models (LLMs) growing capabilities continue to reshape NLP, enabling impressive generalization across diverse tasks, including instruction following, reasoning, summarization, translation, and tool use.
 With the advent of general-purpose foundation models such as GPT-4 (achiam2023gpt), Claude (anthropic2024claude35sonnet), Gemini (team2023gemini), Deepseek (liu2024deepseek), and Llama-3 (touvron2023llama), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems.
@@
 Core Flexibility Advanced Capabilities Usability & Specialization
-Framework
-Custom
-Data
-Custom
-Models
-Custom
-Prompt
-Quant-
-ization
-Long
-Context
-Tool
-Use
-Dist.
-Infer.
-Visual
-Analysis
-Zero-Code
-UI
-Interact.
-CLI
-Low-Resource
-Multilingual
+Framework Custom Data Custom Models Custom Prompt Quant- ization Long Context Tool Use Dist. Infer. Visual Analysis Zero-Code UI Interact. CLI Low-Resource Multilingual
 lm-eval-harness
@@
 indic-eval
+UltraEval
+GlotEval
 Eka-Eval
 Table 1: Feature comparison with state-of-the-art frameworks.
-Takeaway: Eka-Eval seamlessly integrates an accessible dual interface (Zero-Code UI + Interactive CLI) and cutting-edge features (Long Context, Tool Use) with Low-Resource Multilingual benchmarks.
-An Alarming Need: Existing evaluation frameworks for LMs, such as HELM (liang2022helm), EleutherAI Harness (gao2021framework), lm-eval-harness (gao2021framework), and OpenCompass (wu2023opencompass), predominantly target high-resource languages and rely on command-line interfaces with limited benchmark coverage (watts2024pariksha).
+Takeaway: Eka-Eval is the only framework to achieve full coverage across all eleven dimensions, uniquely combining a zero-code UI and low-resource multilingual support with advanced capabilities such as long-context and tool-use evaluation.
+An Alarming Need: Existing evaluation frameworks for LMs, such as HELM (liang2022holistic), lm-eval-harness (gao2021framework), OpenCompass (wu2023opencompass), DeepEval (deepeval), and UltraEval (he-etal-2024-ultraeval) predominantly target high-resource languages and rely on command-line interfaces with limited benchmark coverage (watts2024pariksha).
 These frameworks create significant barriers for non-technical users who require graphical interfaces and researchers working with low-resource and multilingual settings.
 The challenge is especially pronounced in linguistically diverse regions globally, where numerous languages serve large populations; yet, existing evaluation frameworks lack comprehensive multilingual support and culturally grounded benchmarks, limiting rigorous assessment of language model capabilities across diverse linguistic settings (pomerenke2025ai; xuan-etal-2025-mmlu; vayani2025all).
-Motivation: Existing evaluation frameworks suffer from several limitations such as (1) absence of user-friendly interfaces, requiring extensive command-line expertise, broken-pipelines and manual configuration (gao2021framework; liang2022holistic), (2) restricted flexibility in benchmark selection and task customization (wang2018glue; srivastava2022beyond), (3) insufficient task diversity, with fragmented support for advanced capabilities like long-context reasoning and tool use (qin2023toolllm; zhang2024bench), and (4) limited multilingual coverage, particularly for low-resource Indian languages (hu2020xtreme; ahuja2023mega).
+Motivation: Existing evaluation frameworks suffer from several limitations, such as (1) absence of user-friendly interfaces, requiring extensive command-line expertise, broken-pipelines, and manual configuration (gao2021framework; liang2022holistic), (2) restricted flexibility in benchmark selection and task customization (wang2018glue; srivastava2022beyond), (3) insufficient task diversity, with fragmented support for advanced capabilities like long-context reasoning and tool use (qin2023toolbench; zhang-etal-2024-bench), and (4) limited multilingual coverage, particularly for low-resource languages (ahuja-etal-2023-mega; luo-etal-2025-gloteval).
 To address this, we propose Eka-Eval, a uniquely scalable and extensible evaluation framework that unifies multilingual benchmarks under flexible task configurations.
@@
 Contributions: The key contributions are:
-  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 50 multilingual benchmarks across nine major evaluation categories.
-  • We introduce a first-ever Zero-Code UI for benchmark evaluation in diverse and low-resource multilingual settings, along with visualisations, leaderboard, and an interactive CLI, making large-scale evaluation accessible to both non-technical and technical users.
+  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 55 multilingual benchmarks across nine major evaluation categories.
+  • We introduce an open-source LLM evaluation framework that unifies a zero-code UI, an interactive CLI, and built-in support for low-resource multilingual settings, along with visualisations, and leaderboard, making large-scale evaluation accessible to both non-technical and technical users.
   • Through a comprehensive human evaluation involving eleven participants comparing five existing frameworks, showing that Eka-Eval achieves the highest participant ratings across all six features, performing at least 2× better than existing frameworks in key usability and effectiveness.
 2 Relevant Work
-General-Purpose Frameworks: Early multi-task suites such as GLUE (wang2018glue) and SuperGLUE (wang2019superglue) established foundational standards for evaluation.
-More comprehensive frameworks, including HELM (liang2022helm), BIG-Bench (srivastava2022beyond), lm-eval-harness (gao2021framework), and OpenAI Evals (openai2023evals) broaden task coverage and introduce extensible evaluation pipelines.
-yet typically require substantial effort and provide limited multilingual capabilities.
-Modular ecosystems such as OpenCompass (wu2023opencompass) and FreeEval (yu2024freeeval) offer distributed inference and advanced features, but still lack unified evaluation and zero-code accessibility.
-Specialized Capability Benchmarks: Several domain-specific benchmarks target advanced LLM abilities, such as tool use in ToolBench (qin2023toolbench), agentic behavior in API-Bank (li2023api) and InfiniteBench (zhang2024bench), long-context reasoning in RULER (hsieh2024ruler), and mathematical and coding competencies in GSM8K (cobbe2021training) and HumanEval (chen2021evaluating).
-While powerful, these benchmarks are fragmented across independent repositories, requiring users to mainitain multiple codebases to conduct a thorough evaluation.
-Multilingual Evaluation: Multilingual suites such as XTREME (hu2020xtreme) and MEGA (ahuja2023mega) cover 30+ languages but often rely on machine-translated content that lacks cultural grounding.
-Recent efforts of MCEval (Huang2025MCEvalAD) (13 cultures, 39K dynamic instances), GlotEval (Luo2025GlotEvalAT) (language-agnostic, culturally adaptive prompts), BLEnD (myung2024blend) (16 cultures, 52K native QA pairs), CulturalBench (chiu2024culturalbench) (45 global regions), and ALM-Bench (vayani2025all) (100 languages with cultural images) address this gap.
+General-Purpose Frameworks: Early multi-task suites such as GLUE (wang2018glue) and SuperGLUE (wang2019superglue) established foundational standards.
+More comprehensive frameworks, including HELM (liang2022holistic), BIG-Bench (srivastava2022beyond), and lm-eval-harness (gao2021framework), introduced extensible evaluation pipelines but typically require substantial technical effort and provide limited multilingual capabilities.
+Modular ecosystems such as FreeEval (yu2024freeeval) and OpenCompass (wu2023opencompass) offer advanced features like distributed inference.
+While OpenCompass covers extensive benchmarks and zero-code UI for limited models, custom evaluations still require CLI interaction or manual YAML configuration.
+Specialized Capability Benchmarks: Several domain-specific benchmarks target advanced LLM abilities, such as tool use in ToolBench (qin2023toolbench), agentic behavior in API-Bank (li2023api) and InfiniteBench (zhang-etal-2024-bench), long-context reasoning in RULER (hsieh2024ruler), and mathematical and coding competencies in GSM8K (cobbe2021training) and HumanEval (chen2021evaluating).
+While powerful, these benchmarks are fragmented across independent repositories, requiring users to maintain multiple codebases to conduct a thorough evaluation.
+Multilingual Evaluation: Multilingual suites such as XTREME (hu2020xtreme) and MEGA (ahuja-etal-2023-mega) cover 30+ languages but often rely on machine-translated content that lacks cultural grounding.
+Recent efforts of MCEval (Huang2025MCEvalAD) (13 cultures, 39K dynamic instances), GlotEval (luo-etal-2025-gloteval) (language-agnostic, culturally adaptive prompts), BLEnD (myung2024blend) (16 cultures, 52K native QA pairs), CulturalBench (chiu2024culturalbench) (45 global regions), and ALM-Bench (vayani2025all) (100 languages with cultural images) address this gap.
 However, these benchmarks still face insufficient data for low-resource languages, cultural biases favor dominant dialects, and evaluation metrics struggle to capture cultural appropriateness.
-Eka-Eval mitigates these gaps by unifying 50+ diverse benchmarks within a single modular workflow.
-It offers dual accessibility (zero-code UI + interactive CLI), advanced capabilities (long-context, tool-use, distributed inference), and multilingual language support for comprehensive evaluation.
+Existing frameworks are often complex and poorly modularized, hindering seamless incorporation into research workflows (he-etal-2024-ultraeval).
+Furthermore, most toolkits also remain disproportionately focused on English and a handful of high-resource languages, making non-English-centric evaluations less practical for many researchers (luo-etal-2025-gloteval), and none offer a zero-code UI or native support for low-resource languages, creating a gap for practitioners evaluating models in low-resource settings without dedicated systems expertise.
+Eka-Eval addresses these gaps by unifying 55+ benchmarks across nine evaluation categories within a single modular workflow, with a strong focus on low-resource multilingual evaluation.
+It is, to our knowledge, the only framework to pair a zero-code web UI along with interactive CLI, easy-to-use framework making rigorous evaluation accessible to non-technical users.
 3 Capabilities of LLM Evaluation Frameworks
@@
 (7) Distributed Inference: Parallelized evaluation across multiple compute nodes.
-(8) Visual Analysis: Generation of interpretable visualizations (bar charts, radar plots, heatmaps, etc.).
+(8) Visual Analysis: Generation of interpretable visualizations (bar charts, plots, heatmaps, etc.).
 C. Usability & Specialization:
-(9) Zero-Code UI: Web-based interface enabling evaluation without programming.
+(9) Zero-Code UI: Web-based interface enabling end-to-end evaluation workflows without coding.
 (10) Interactive CLI: Command-line interface for configuring datasets, models, and strategies.
 (11) Low-Resource Multilingual: Comprehensive multilingual benchmarks featuring language-specific tokenization.
-Table 1 presents a feature-level comparison demonstrating that Eka-Eval is the only framework offering complete coverage across technical, linguistic, and usability dimensions.
-While lm-eval-harness, OpenCompass, and HELM provide strong technical capabilities, they require substantial expertise and suffer from usability challenges.
-indic-eval, though tailored to Indic languages, supports only seven benchmarks and lacks broader functionality.
-Critically, existing frameworks lack zero-code web interfaces, restricting accessibility to users with programming expertise.
-Eka-Eval uniquely combines comprehensive evaluation capabilities with an intuitive zero-code UI, significantly lowering barriers to rigorous multilingual assessment.
+Table 1 demonstrates that Eka-Eval achieves complete technical, linguistic, and usability coverage.
+While lm-eval-harness and HELM offer strong capabilities, they demand substantial programming expertise.
+OpenCompass includes a zero-code UI, but restricts it to leaderboard models and pre-configured benchmarks.
+OpenAI Evals and DeepEval support custom pipelines but lack quantization, long-context, and multilingual capabilities.
+Similarly, UltraEval and GlotEval provide provide robust customization and visual analysis (with the latter targeting low-resource languages) but but lack tool-use, long-context evaluation, and a zero-code UI.
+indic-eval is limited to seven benchmarks and offers no general-purpose capabilities.
+Critically, no existing framework pairs a zero-code UI with native low-resource multilingual support; Eka-Eval addresses both, lowering barriers to rigorous multilingual assessment.
 4 Eka-Eval
@@
 It provides a zero-code, web-based interface that allows users to run benchmarks, adjust parameters, and visualize results through interactive analytics and leaderboards, while a CLI enables detailed configuration and large-scale evaluations.
-The framework supports flexible model selection for both local and API-based models and is built around three core principles: modularity for seamless extension and customization; accessibility to support diverse environments, including low-resource language settings; and comprehensiveness to cover a wide range of capabilities and multilingual benchmarks, with particular attention to underserved areas such as long-context reasoning and tool use.
+The framework supports flexible model selection for both local and API-based models and is built around three core principles: modularity for seamless extension and customization; accessibility to support diverse environments, including low-resource language settings; and comprehensiveness to cover a wide range of capabilities and low-resource multilingual benchmarks, with particular attention to underserved areas such as long-context reasoning and tool use.
 4.2 System Architecture
-The framework has a layered architecture with four main components (illustrated in Figure 1):
+The framework has a layered architecture with four main components (as illustrated in Figure 1):
 4.2.1 Evaluation Engine
@@
 Table 2: Average participant ratings of the evaluation frameworks by eleven participants (mean $\pm$ standard deviation; Likert scale: 1–5, 1 = poor and 5 = excellent).
-Takeaway: Eka-Eval achieves the highest ratings across all functionalities, with participants consistently perceiving it as superior across the six assessed capabilities.
+Takeaway: Eka-Eval achieves the highest mean ratings across all six dimensions, with notably stronger scores on Zero-Code UI and Global Low-Resource Language Support.
    Framework    Time Taken
@@
    Eka-Eval    $\mathbf{11\pm 3.18}$
-Table 3: Comparison of time taken for installation and configuration across six frameworks reported across five successful runs on GPU.
-Takeaway: Eka-Eval records the lowest setup time and exhibits consistently low variance.
+Table 3: Comparison of time taken for installation and configuration across six frameworks reported across five successful runs on GPU-setup.
+Takeaway: Eka-Eval records the lowest setup time and variance.
 4.2.4 Results Processing System
 Handles comprehensive output management through three secondary components:
-Metrics Calculator: Employs HuggingFace’s industry-standard evaluate library to compute essential metrics such as accuracy, BLEU, F1-score, exact match, and Pass@1, ensuring consistency with widely adopted benchmarks and standard evaluation practices across diverse LLM tasks.
+Metrics Calculator: Employs HuggingFace’s industry-standard evaluate library to compute essential metrics such as accuracy, BLEU (papineni2002bleu), F1-score (10.1145/3606367), exact match (rajpurkar2016squad), and Pass@1 (chen2021evaluating), ensuring consistency with widely adopted benchmarks and standard evaluation practices across diverse LLM tasks.
 Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps, and radar plots (including support for cross-model comparisons).
@@
 Handles complex evaluation logic with ease of use through a streamlined full-stack user interface (See Appendix §B and Figure 4).
-Full-Stack Architecture: A decoupled React frontend and FastAPI backend, served through Nginx, provide a stable and scalable deployment layer.
-Real-Time Telemetry: WebSocket-based streaming delivers live inference logs and GPU status directly to the UI, eliminating the need for CLI monitoring.
+Full-Stack Architecture: A decoupled React fedosejev2015react frontend and FastAPI tiangolo2018fastapi backend, served through Nginx, provide a stable and scalable deployment layer.
+Real-Time Telemetry: WebSocket-based streaming delivers live inference logs and GPU status directly to the UI, eliminating the need for CLI.
 LLM-Powered Diagnostics: An LLM (LLaMA-3.3 70B (grattafiori2024llama)) interprets evaluation outputs and generates summaries of model performance and failure patterns, revealing additional insights.
-Interactive Dashboard: A “Zero-Code” interface with visual comparisons and leaderboards enables users to run and analyze evaluations without technical overhead.
+Interactive Dashboard: A “Zero-Code” user-interface with visual comparisons and live leaderboards enables users to run and analyze evaluations without technical expertise.
 4.3 Comprehensive Benchmark Coverage
-Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 50+ benchmarks (See Appendix A).
+Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 55+ benchmarks (See Appendix A).
 These categories include:
   1. 1.
-    Code Generation and Programming: Programming abilities are assessed using HumanEval (chen2021evaluating), MBPP (austin2021program), HumanEval+ (liu2023your), and MBPP EvalPlus (liu2023your) with Pass@1 metrics.
+    Code Generation and Programming: These are assessed using HumanEval (chen2021evaluating), MBPP (austin2021program), HumanEval+ (liu2023your), PythonSaga (yadav-etal-2024-pythonsaga) and MBPP EvalPlus (liu2023your) with Pass@1 metrics.
   2. 2.
@@
   3. 3.
-    Reading Comprehension: Text understanding is evaluated using SQuAD (rajpurkar2018know), QuAC (choi2018quac), and BoolQ (clark2019boolq) with F1 and exact match metrics.
+    Reading Comprehension: Text understanding is evaluated using SQuAD (rajpurkar2018know), QuAC (choi-etal-2018-quac), and BoolQ (clark2019boolq) with F1 and exact match metrics.
   4. 4.
@@
   6. 6.
-    Long-Context Understanding: For extended context, we include ZeroSCROLLS (shaham2023zeroscrolls), Needle-in-a-Haystack (wang2024multimodal), and InfiniteBench (zhang2024bench).
+    Long-Context Understanding: For extended context, we include ZeroSCROLLS (shaham2023zeroscrolls), Needle-in-a-Haystack (wang-etal-2025-multimodal), and InfiniteBench (zhang-etal-2024-bench).
   7. 7.
-    General Reasoning: Foundational capabilities are tested via MMLU (hendrycks2021mmlu), MMLU-Pro (wang2024mmlu), IFEval (zhou2023instruction), BBH (suzgun2022challenging), and AGI-Eval (zhong2023agieval).
+    General Reasoning: Foundational capabilities are tested via MMLU (hendrycks2021mmlu), MMLU-Pro (wang2024mmlu), IFEval (zhou2023instruction), BBH (suzgun2022challenging), and AGI-Eval (zhong-etal-2024-agieval).
   8. 8.
@@
   9. 9.
-    Multilingual Language Support: Eka-Eval features the largest unified Multilingual suite (20 benchmarks), including: (i) Knowledge: IndicMMLU-Pro (sankalp2025indicmmlupro), MMLU-IN (hendrycks2020measuring), TriviaQA-IN (joshi2017triviaqa), MILU (verma2024milu).
-(ii) Reasoning: HellaSwag-IN (zellers2019hellaswag), ARC-C IN (clark2018think), IndicCOPA (kakwani2020indicnlpsuite), GSM8K-IN (cobbe2021training).
-(iii) Reading & QA: BoolQ-IN (clark2019boolq), XQuAD-IN (artetxe2019cross), XorQA-IN (asai2021xor), Indic-QA (singh2025indicqabenchmarkmultilingual).
-(iv) Generation (NLG): Flores-IN (goyal-etal-2022-flores), IndicParaphrase, IndicWikiBio, IndicQuestionGeneration, IndicSentenceSummarization, IndicHeadlineGeneration (singh2024indicgenbench).
-(v) NLU: IndicNER, IndicSentiment, IndicGLUE (kakwani2020indicnlpsuite).
+    Multilingual and Low-Resource Language Support: LLMs are known to underperform on languages spoken across South Asia, Africa, and Southeast Asia, yet no existing evaluation framework provides sufficient coverage to diagnose this gap (luo-etal-2025-gloteval; kakwani2020indicnlpsuite).
+Eka-Eval features the largest unified Multilingual suite (23 benchmarks), including: (i) Knowledge: IndicMMLU-Pro (sankalp2025indicmmlupro), MMLU-IN (hendrycks2021measuring), TriviaQA-IN (joshi2017triviaqa), MILU (verma-etal-2025-milu) spanning Kannada, Odia, Malayalam, and Urdu.
+(ii) Reasoning: HellaSwag-IN (zellers2019hellaswag), ARC-C-IN (clark2018think), IndicCOPA (kakwani2020indicnlpsuite), XCOPA (ponti-etal-2020-xcopa) in Quechua, Indonesian, and Swahili, and GSM8K-IN (cobbe2021training) for reasoning across Indic languages.
+(iii) Reading & QA: Belebele (bandarkar-etal-2024-belebele) covering 122 languages including Yoruba, Oromo, and Indonesian; BoolQ-IN (clark2019boolq), XQuAD-IN (artetxe2020cross) in Hindi and Greek, XorQA-IN (asai2021xor) in Bengali and Telugu, and Indic-QA (singh2025indicqabenchmarkmultilingual).
+(iv) Generation (NLG): Flores-IN (goyal-etal-2022-flores) for translation across 200 languages; IndicParaphrase, IndicWikiBio, IndicQuestionGeneration, IndicSentenceSummarization, and IndicHeadlineGeneration (singh2024indicgenbench) in Odia, Assamese and Punjabi.
+(v) NLU: IndicNER and IndicSentiment in 11 languages, IndicGLUE (kakwani2020indicnlpsuite), and XNLI (conneau-etal-2018-xnli) for cross-lingual inference in Arabic, Swahili, and Urdu.
+5 Experiments
+To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval.
+These frameworks were selected as representative state-of-the-art baselines spanning across widely adopted general-purpose to specialized multilingual evaluation suites, ensuring a rigorous comparison.
+Evaluation Procedure
+Eleven graduate-level computer science researchers were selected from NLP and software engineering programs.
+Participants were stratified by prior experience using at least one existing evaluation framework (5 with, 6 without) and evaluated each framework in a single-blind setup.
+Selection was not tied to knowledge of Eka-Eval, mitigating demand bias; sample size was constrained by available GPU resources.
+Participants followed a standardized evaluation protocol in a controlled environment.
+They installed each framework, integrated HuggingFace models - including gemma-2-2b (team2024gemma) and sarvam-1B sarvamai_sarvam1_2024, and executed the recommended workflows on four diverse benchmarks: WinoGrande, PIQA, ARC-C-IN and MMLU-IN.
+All ratings were recorded using the standardized protocol (Appendix §D) to ensure direct and consistent comparison across frameworks.
+Participants assessed each framework across six evaluation criteria and rated them using a Likert scale ranging from 1 to 5:
+  1. 1.
+    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and run an initial benchmark.
+  2. 2.
+    Ease of Navigation: Intuitiveness of navigation, benchmark selection, and configuration, including CLI clarity, documentation quality, and ease of discovering options.
+  3. 3.
+    Zero-Code UI Availability: Presence and usability of GUI-based tools for visualization, benchmark execution, and result analysis, eliminating the need for command-line interaction.
+  4. 4.
+    Result Reporting and Export: Clarity and accessibility of evaluation outputs, with options to export results (e.g., JSON, CSV) and create visualizations such as bar charts or heatmaps.
+  5. 5.
+    Extensibility: Ease of customizing the framework to add new prompt templates, models, benchmarks, or evaluation metrics.
+  6. 6.
+    Multilingual Language Support: Support for multilingual benchmarks such as Belebele, ARC-IN, FLORES, and XNLI.
+Table 2 presents average participant ratings across six evaluation criteria using a 1-5 Likert scale (1= poor or absence of functionality, 5= excellent or flexible capability).
+Eka-Eval achieved the highest ratings across all criteria, establishing superior usability and capability.
+Moreover, Table 3 summarizes the average setup times across six frameworks, which varied widely due to dependency issues and unstable codebases in several baselines.
+Nine of the eleven participants reported that HELM suffered from poor documentation and a broken pipeline, while OpenCompass required manual code edits and indic-eval experienced installation difficulties.
+Eka-Eval achieved the shortest setup time.
+Table 4 (See Appendix 4 for full results) presents reproduced benchmark scores on gemma-2-2b, showing how they align with broader framework differences and highlighting Eka-Eval’s consistency across global-multilingual benchmarks.
+HELM was excluded from this comparison, as it neither supports these benchmarks nor produced functional results despite repeated attempts.
+6 Conclusion and Future Work
+We present Eka-Eval, an open-source unified evaluation framework designed to overcome persistent challenges in accessibility, reproducibility, and low-resource multilingual coverage for language model assessment.
+Our study with eleven participants demonstrates substantial usability improvements, achieving lowest setup time and the highest satisfaction ratings.
+Eka-Eval enables an interactive UI, rigorous, reproducible evaluation without manual coding expertise.
+The current release is an ongoing work, with 30 stars and 3 forks on GitHub.
+Future work will expand coverage to multimodal LLMs, benchmarks, and AI diagnostic capabilities.
+7 Limitations
+Multimodal Support.
+Eka-Eval currently supports text-only LLMs; multimodal model evaluation is left to future work.
+Reproducibility.
+Evaluation results may be affected by changes in external dataset or model versions; we recommend explicit versioning and caching for fully reproducible runs.
+User study scope.
+The usability study (n=11) constitutes a formative evaluation conducted under GPU resource constraints, providing initial usability insights and motivating broader future evaluations.
+Inference Backend.
+Eka-Eval currently relies on HuggingFace Transformers for model inference; vLLM backend integration is underway to enable faster, high-throughput evaluation at scale.
+Scalability.
+The current release processes requests sequentially; supporting concurrent multi-user deployments may require additional infrastructure and compute beyond the scope of this work.
+8 Ethics Statement
+This research uses publicly available data without personally identifiable information.
+All datasets and models comply with their terms of use.
+The work is intended for academic research.
+Potential misuse or unintended amplification of biases should be carefully considered before deployment.
+9 Acknowledgements
+The authors express their gratitude to Himanshu Beniwal for helping with writing the manuscript and to Birudugadda Srivibhav, Sailesh Panda, Gautham Bharati, Indrayudh Mandal, Krudant Randai, Sahil Gawande, Tirth Bhatt, Shruti Bhat, Mann agarwal, Himanshu Sharma and Khushbu Bijawat for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results.
+We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Mahavir Patil.
+Appendix A Appendix
    Framework    Winogrande    PIQA    ARC-IN    MMLU-IN
@@
 Takeaway: Eka-Eval reproduces verified consistent scores with established frameworks and closely aligns with their actual benchmark results.
-4.4 Extensibility and Customization
-Eka-Eval emphasizes extensibility through a low-code plugin architecture that enables easy integration of new benchmarks and custom metrics with minimal modifications, managed via a hierarchical JSON configuration system that supports complex setups such as parameter sweeps and prompt variations.
-The framework provides a zero-code UI that enables real-time, interactive adjustment of inference parameters (e.g., temperature, batch size) and evaluation strategies, eliminating the need for users to manage configuration files directly.
-5 Experiments
-To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval.
-Evaluation Procedure
-Eleven graduate-level computer science researchers with demonstrated proficiency in Git/GitHub evaluated each framework in a single-blind setup.
-Participants followed a standardized evaluation protocol in a controlled environment.
-They installed each framework, integrated HuggingFace models - including gemma-2-2b (team2024gemma) and sarvam-1B sarvamai_sarvam1_2024, and executed the recommended workflows on four diverse benchmarks: WinoGrande, PIQA, ARC-C-IN and MMLU-IN.
-They then rated the framework on five criteria using a Likert scale ranging from 1 to 5:
-EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks.
-Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility, with explicit dos and don’ts specified for all participants.
-They were permitted to consult official documentation but were prohibited from seeking external technical support beyond publicly available framework resources.
-Participants assessed each framework across six evaluation criteria:
-  1. 1.
-    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and run an initial benchmark.
-  2. 2.
-    Ease of Navigation: Intuitiveness of navigation, benchmark selection, and configuration, including CLI clarity, documentation quality, and ease of discovering options.
-  3. 3.
-    Zero-Code UI Availability: Presence and usability of GUI-based tools for visualization, benchmark execution, and result analysis, eliminating the need for command-line interaction.
-  4. 4.
-    Result Reporting and Export: Clarity and accessibility of evaluation outputs, with options to export results (e.g., JSON, CSV) and create visualizations such as bar charts or heatmaps.
-  5. 5.
-    Extensibility: Ease of customizing the framework to add new prompt templates, models, datasets, or evaluation metrics.
-  6. 6.
-    Multilingual Language Support: Support for multilingual benchmarks such as ARC-IN, FLORES, and similar. Evaluators assessed the availability of prebuilt resources.
-Table 2 presents average participant ratings across six evaluation criteria using a 1-5 Likert scale (1= poor or absence of functionality, 5 =excellent or flexible capability).
-Eka-Eval achieved the highest ratings across all criteria, establishing superior usability and capability.
-Moreover, Table 3 summarizes the average setup times across six frameworks, which varied widely due to dependency issues and unstable codebases in several baselines.
-Nine of the eleven participants reported that HELM suffered from poor documentation and a broken pipeline, while OpenCompass required manual code edits and indic-eval experienced installation difficulties.
-Eka-Eval achieved the shortest setup time.
-Finally, Table 4 presents reproduced benchmark results on gemma-2-2b, showing how they align with broader framework differences and highlighting Eka-Eval’s consistency across global-multilingual benchmarks.
-HELM was excluded from this comparison, as it neither supports these benchmarks nor produced functional results despite repeated attempts.
-6 Conclusion and Future Work
-We present Eka-Eval, an open-source unified evaluation framework designed to overcome persistent challenges in accessibility, reproducibility, and multilingual coverage for language model assessment.
-Our study with eleven participants demonstrates substantial usability improvements, achieving lowest setup time and the highest satisfaction ratings.
-Eka-Eval enables an interactive UI, rigorous, reproducible evaluation without manual integration overhead.
-The current release is an ongoing work, with 30 stars and 3 forks on GitHub.
-Future development will extend the framework to more than 100 tasks, specialized domains, and advanced diagnostic capabilities, establishing Eka-Eval 2.0 as a comprehensive and equitable standard for multilingual language model evaluation.
-7 Limitations
-While Eka-Eval supports a wide range of benchmarks and model backends, it does not yet support multi-modal LLMs.
-Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
-8 Ethics Statement
-This research uses publicly available data without personally identifiable information.
-All datasets and models comply with their terms of use.
-The work is intended for academic research.
-Potential misuse or unintended amplification of biases should be carefully considered before deployment.
-9 Acknowledgements
-The authors express their gratitude to Himanshu Beniwal for helping with writing the manuscript and to Birudugadda Srivibhav, Sailesh Panda, Gautham Bharati, Indrayudh Mandal, Krudant Randai, Sahil Gawande, Tirth Bhatt, Shruti Bhat, Mann agarwal, Himanshu Sharma and Khushbu Bijawat for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results.
-We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Mahavir Patil.
-Appendix A Appendix
 A.1 Multilingual benchmark configuration
@@
 Users can customize prompt strategies and easily configure them in the benchmark configuration file, as shown in Figure 2.
-Figure 4: High-level architectural overview of the Eka-Eval framework.
-Appendix B Zero-code Web Interface
-To democratize access for LLM evaluation, Eka-Eval includes a comprehensive "Zero-code" web interface.
-This GUI mirrors the capabilities of the CLI serving as a frontend layer as seen in (Figure 4) while adding advanced features for prompt configuration, resource management, and AI-driven analysis making it accessible for both non-technical and technical users.
-Figure 5: High-level benchmark selection dashboard categorized by domain (e.g., Math, Code Generation, Indic Benchmarks).
+Figure 4: The overall architectural pipeline of the Eka-Eval framework.
+Appendix B Zero-code User Interface
+To democratize access for LLM evaluation, Eka-Eval includes a “Zero-code” user interface that mirrors the full capabilities of the CLI while adding advanced features for prompt configuration, resource management, and AI-assisted analysis, making rigorous evaluation accessible to both non-technical and technical users.
+Figure 4 illustrates the end-to-end architecture, spanning user input, model inference, evaluation execution, and result visualization across Eka-Eval’s four core components.
+Figure 5: Standardized study protocol specifying permitted and prohibited actions for all participants.
+Figure 6: Participants assessment form.
+Figure 7: ModelHub selection panel to choose any HuggingFace/Local models for benchmark evaluation.
+Figure 8: Benchmark selection dashboard categorized by domain (e.g., Math, Code Generation, and other benchmarks.
+Figure 9: Advanced configuration panel for adjusting inference parameters (temperature, batch size) and managing GPU resources.
+Figure 10: The prompt customization interface, allowing users to modify templates and injection points for specific tasks.
+Figure 11: Post-evaluation summary showing overall progress, specific benchmark scores, and options to export results or view detailed analysis.
+Figure 12: AI diagnosis dashboard providing automated textual insights into model strengths and weaknesses based on performance.
+Figure 13: AI diagnosis showing the wrong result of selected benchmark
+Figure 14: Model Leaderboard through aggregated average scores across multiple benchmarks and through live evaluation
+Figure 15: Model Leaderboard featuring multi-dimensional visualizations (radar and bar charts) for comparative analysis.
+Figure 16: Model Key strengths and weaknesses featuring multi-dimensional visualizations along with benchmarks performance analysis for improvement.
 The Benchmark Selection Dashboard
 (Figure 8) allows users to build their own evaluation suite by simply toggling and selecting the benchmarks available and also adding custom benchmarks.
-Figure 6: Advanced configuration panel for adjusting inference parameters (temperature, batch size) and managing GPU resources.
-For advanced control over the evaluation process, Configuration Panel (Figure 6) offers adjustable settings such as temperature, top-p and batch size.It also includes a built-in GPU resource manager, allowing users to allocate specific or multi CUDA for evaluation.
-Figure 7: The prompt customization interface, allowing users to modify templates and injection points for specific tasks.
-A key advantage of the web interface is the Prompt Customization Engine (Figure 7).
-It lets users visually edit system prompts, few-shot templates, and variable fields, removing the need to manually modify JSON configuration files.
-Figure 8: Post-evaluation summary showing overall progress, specific benchmark scores, and options to export results or view detailed analysis.
+For advanced control over the evaluation process, Configuration Panel (as shown in Figure 9) offers adjustable settings such as temperature, top-p and batch size.
+It also includes a built-in GPU resource manager, allowing users to allocate specific or multi-GPU support for evaluation.
+A key advantage of the web user interface is the Prompt Customization Engine (Figure 10).
+It lets users visually edit system prompts, few-shot templates, and variable fields, GPU allocation, thus removing the need to manually modify JSON configuration files.
 The Evaluation Dashboard
-(Figure 8) shows real time progress during inference including a live console and, once finished, provides a summary of the model’s performance with options to download logs or explore detailed analysis.
-Figure 9: AI diagonosis dashboard providing automated textual insights into model strengths and weaknesses based on performance.
+Figure 11 shows real time progress during inference including a live console and, once finished, it provides a summary of the model’s performance with options to download logs or explore detailed analysis.
 AI Diagnosis Dashboard
-uses LLama 3.3-70b to analyze the evaluation logs, generating summaries of the model’s observed strengths (e.g., "Strong logical coherence") and weaknesses (e.g., "Hallucination in low-resource languages") as example shown in (Figure 9).
-Figure 10: AI diagnosis of the wrong result of benchmark selected
+uses LLama 3.3-70b model to analyze the evaluation logs, generating summaries of the model’s observed strengths (e.g., ‘Strong logical coherence’) and weaknesses (e.g., ‘Hallucination in low-resource languages’) as example shown in Figure 12.
 Users can also inspect specific failure cases (Figure 13).
 Selecting a failed benchmark item shows the prompt, the model’s wrong answer, and an AI-generated explanation of why the error likely occurred, making it easy to spot and fix issues.
-Figure 11: Model Leaderboard featuring multi-dimensional visualizations (radar and bar charts) for comparative analysis.
-Finally, the Model Leaderboard (Figure 11) provides a central place to compare models.
-It gathers results from all local runs and visualizes performance differences with radar charts and bar graphs, helping users choose the best model.
+Finally, the Model Leaderboard Figure 14, Figure 15 and Figure 16 provides a centralized, dynamic interface for comparative analysis.
+Unlike static evaluation tables, it aggregates results from all local runs securely.
+Researchers can filter outputs by model size, language, or specific benchmark, and generate side-by-side multi-dimensional radar charts.
+To maintain integrity, the leaderboard includes version tracking for datasets and prevents score manipulation by linking directly to the immutable evaluation logs generated by the backend engine.
 Appendix C Interactive CLI
@@
 Figure 18: Model selection in the Eka-Eval framework.
-Figure 14: Interactive visualisation setup in Eka-Eval.
+Figure 19: Interactive and user-friendly visualisation setup in Eka-Eval.
 Figure 20: Subtask selection within a task group.
-Figure 16: Bar chart visualisation
+Figure 21: Bar chart visualisation on Reading Comprehension benchmarks
 After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain (see Figure 20).
@@
 Figure 21 shows performance breakdown across sub-tasks like BoolQ, SQuAD, and QuAC. It provides intuitive insight into strengths and weaknesses of the model.
-Appendix D Computation Requirement and Budget
+Appendix D User Study Protocol
+Participants received a standardized evaluation protocol and rating form (Figures 5 and 6) to ensure consistency across all framework evaluations.
+Each participant was permitted to consult official documentation but was prohibited from seeking external technical assistance beyond publicly available framework resources.
+Explicit guidelines specified permitted and prohibited actions to minimize evaluation bias.
+Appendix E Extensibility and Customization
+Eka-Eval emphasizes extensibility through a low-code plugin architecture that enables easy integration of new benchmarks and custom metrics with minimal modifications, managed via a hierarchical JSON configuration system that supports complex setups such as parameter sweeps and prompt variations.
+The framework provides a zero-code UI that enables real-time, interactive adjustment of inference parameters (e.g., temperature, batch size) and evaluation strategies, eliminating the need for users to edit the configuration files directly.
+Appendix F Computation Requirement and Budget
 For computational infrastructure, experiments were carried out on four NVIDIA Tesla V100 32 GB GPUs, with an estimated cost of \$5,431.20 per month based on Google Cloud Platform (GCP)  [^1] https://cloud.google.com/products/calculator Calculator pricing.